### v1

# Stacky Batyrev–Manin conjecture and modular curves

Abstract.
Let $\mathscr{X}_{0}(N)$ be the Deligne–Rapoport modular stack of elliptic curves endowed with a cyclic rational $N$-isogeny over a number field $F$. Let $N\in\{1,2,3,4,5,6,7,8,9,10,12,13,16,18,25\},$ which are precisely the values for which the coarse moduli space of $\mathscr{X}_{0}(N)$ is isomorphic to $\mathbb{P}^{1}$. We show that the stacky Batyrev–Manin conjecture [DY24] holds for the naive height on $\mathscr{X}_{0}(N)$ when $F=\mathbb{Q}$. In the process, we give a concrete description of $\mathscr{X}_{0}(N)$ as a square root stack over a stacky curve.
Key words and phrases:
Batyrev-Manin conjecture, modular curves, Deligne-Mumford stacks, orbifold pseudoeffective cone
2020 Mathematics Subject Classification:
Primary 11G50, 14G05; Secondary 14A20, 11G18
1. Introduction
1.1. Batyrev–Manin conjecture and stacks
1.1.1.
The standard Batyrev–Manin conjecture predicts the asymptotic behavior of the number of rational points of bounded height on rationally connected varieties. It is conjectured in [Pey95, BT98] that the number of points of height bounded by $B$ grows as $CB^{a}\log(B)^{b-1}$ for certain constants $C>0$, $a>0$, and $b\geq 1$. The constants $a$ and $b$, which are the main focus of this paper, are expressed in terms of relative positions of the line bundle defining height and the canonical line bundle to the effective cone of the variety.
Recently, the Batyrev–Manin conjecture has been studied for Deligne–Mumford stacks (short DM stacks) [ESZB23, DY24, DY25b]. The motivation comes from the fact that some counting results and predictions, such as Malle’s conjecture [Mal04] or such as counting elliptic curves, with or without a level structure, of bounded height [BS24, Mol23, Phi22, MV23], often feature the formula of the same shape $CB^{a}\log(B)^{b-1}$. They turn out to be concerned with rational points of DM stacks. The main conjecture of [DY24] gives a prediction for the number of rational points of bounded height for (a reasonable class of) DM stacks.
1.1.2.
A difficulty when formulating the Batyrev–Manin conjecture for stacks is the definition of heights. Namely, a line bundle on a Deligne–Mumford stack does not produce a useful height if some of its powers are pullbacks of a line bundle from the coarse moduli space; in this case, the height is (up to powers) pullback of a height from the coarse moduli spaces, which often fail the Northcott property. In [DY24], the problem is resolved by the use of sectors. Roughly, a sector is a deformation equivalent pair of a geometric point of DM stack and an injection of a finite cyclotomic group schemes $\mu_{k}$ (with $k\geq 1$) into the automorphism group scheme $\operatorname{Aut}(x)$.
To state those precisely, let $\mathcal{X}\not\cong\operatorname{Spec}(F)$ be a geometrically irreducible separated smooth Deligne–Mumford $F$-stack, where $F$ is a fixed number field. We also assume that the coarse moduli space $X$ of $\mathcal{X}$ is a projective geometrically rationally connected variety. A sector of $\mathcal{X}$ is an equivalence class of a pair $(x,g)$ with $x\in\mathcal{X}(\overline{F})$ and $g:\mu_{k}\hookrightarrow\operatorname{Aut}(x)$ and $k\geq 1$, where $(x_{1},g_{1}:\mu_{k}\hookrightarrow\operatorname{Aut}(x_{1}))$ and $(x_{2},g_{2}:\mu_{k}\hookrightarrow\operatorname{Aut}(x))$ are equivalent if there exists a connected $F$-scheme $S$, two geometric points $y_{1},y_{2}:\operatorname{Spec}(\overline{F})\to S$, a morphism $y:S\to\mathcal{X}$ and an injection $h:(\mu_{k})_{S}\to\operatorname{Aut}(x)$ and isomorphisms of the restriction $(y,h)$ at $y_{i}$ with $(x_{i},g_{i})$ for $i=1,2$. The set of sectors is finite [DY24, Page 6]. The untwisted sector is the equivalence class of $(x,\mu_{1}\hookrightarrow\operatorname{Aut}(x))$ for a point $x\in\mathcal{X}(\overline{F})$. The other sectors are called twisted sectors.
Using the notion of sectors, the first author and Yasuda in [DY24, Definition 4.3] defined a height of $\mathcal{X}$ from a pair of a $\mathbb{Q}$-line bundle and a function defined on the set of twisted sectors. The rational Néron–Severi space is replaced by the space
$\operatorname{NS}_{\operatorname{orb}}(\mathcal{X}):=\operatorname{NS}(\mathcal{X})_{\mathbb{Q}}\bigoplus_{y\in\pi_{0}^{*}(\mathcal{J}_{0}(\mathcal{X}))}\mathbb{Q}[y],$
where $\pi_{0}^{*}(\mathcal{J}_{0}(\mathcal{X}))$ is the set of twisted sectors of $\mathcal{X}$. The canonical line bundle is replaced by the element $K_{\mathcal{X},\operatorname{orb}}:=(K_{\mathcal{X}},(\operatorname{age}(y)-1)_{y\in\pi_{0}^{*}(\mathcal{J}_{0}(\mathcal{X}))})\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, where $K_{\mathcal{X}}$ is the canonical line bundle of $\mathcal{X}$ and $\operatorname{age}$ is the function defined in [DY24, Definition 2.23]. Finally, the effective cone of $\mathcal{X}$ is replaced by the orbifold pseudoeffective cone, which is explained below.
Let us first define this cone. A stacky curve on $\mathcal{X}_{\overline{F}}$ is a representable $\overline{F}$-morphism $f:\mathcal{C}\to\mathcal{X}_{\overline{F}}$ with $\mathcal{C}$ stacky curve (which is a smooth separated irreducible stack of dimension $1$). By [GS17, Theorem 1] we can assume that the automorphism group schemes of points of $\mathcal{C}$ are $\mu_{k}$ for some $k$. At each stacky point $z$ of $\mathcal{C}$, we get an associated sector $\psi(z)$ of $\mathcal{X}$, it is the one defined by the pair $(f(z),\operatorname{Aut}(z)\hookrightarrow\operatorname{Aut}(f(z))).$ For $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines an intersection number $(f:\mathcal{C}\to\mathcal{X}_{\overline{F}},(L,x))$ by setting it to be the sum of $\deg(f^{*}L)$ and $\sum_{z\in S(\mathcal{C})}x_{\psi(z)}$, where $S(\mathcal{C})$ is the set of stacky points of $\mathcal{C}$ and $x_{0}$ is interpreted as 0. A family of movable stacky curves on $\mathcal{X}_{\overline{F}}$ is a family of stacky curves on $\mathcal{X}_{\overline{F}}$ which dominates $\mathcal{X}_{\overline{F}}$. One defines an intersection number between a family of movable curves and $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, as the generic value of the intersection of the members of the family with $(L,x)$ [DY24, Definition 8.7]. Motivated by the fact that the effective cone is the dual of the cone of movable curves [BDPP13], one defines $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})$ to be the subset of $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$ which intersect non-negatively every movable family of stacky curves on $\mathcal{X}_{\overline{F}}$. If a height is defined by $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines
$a(L,x):=\inf\{t\in\mathbb{R}|\hskip 2.84544ptt\cdot(L,x)+K_{\mathcal{X},\operatorname{orb}}\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})\}$
and
$b(L,x):=\text{codimension of minimal face of }\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})\text{ containing }a(L,x)\cdot(L,x)+K_{\mathcal{X},\operatorname{orb}}.$
Finally, we need one more definition before stating the main conjecture: given a stack $\mathcal{Y}$ and a ring $R$, denote by $\mathcal{Y}\langle R\rangle$ the set of isomorphism classes of $R$-points of $\mathcal{Y}$.
Conjecture 1.1 ([DY24, Conjecture 9.6]).
Assume that $\mathcal{X}\langle F\rangle$ is Zariski-dense in $\mathcal{X}$. Let $(L,x)\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})$ and let $H:\mathcal{X}\langle F\rangle\to\mathbb{R}_{>0}$ be the corresponding height. There exists a thin set $T\subset\mathcal{X}\langle F\rangle$ and $C>0$ such that
$\{\mathcal{X}\langle F\rangle-T|\hskip 2.84544ptH(x)\leq B\}\sim_{B\to\infty}CB^{a(L,x)}\log(B)^{b(L,x)-1}.$
We note two differences with [DY24, Conjecture 9.10]. First, we do not conjecture that the set $T$ is a union of images of breaking thin morphisms [DY24, Definition 9.8]. Second, we do not assume $(L,x)$ is adequate in the sense of [DY24, Definition 5.2]. The condition is relevant for [DY24, Conjecture 5.6], but seems redundant for Conjecture 9.10. (It makes $a$-invariant equal to $1$. In the sequel [DY25b], the adequacy condition is no longer imposed.) Some of the cases we treat in this paper are non-adequate.
1.2. Results
The main theorem of this paper proves the Conjecture 1.1 on several modular curves. Recall that given a stack $\mathcal{X}$ defined over a number field $F$, $\mathcal{X}\langle F\rangle$ is the set of isomorphism classes of $F$-points of $\mathcal{X}$. For each $N\in\mathbb{Z}_{>0}$, let $\mathscr{X}_{0}(N)$ be the compactified Deligne–Rapoport modular stack of elliptic curves endowed with a cyclic rational $N$-isogeny over $F$ (see Definition 3.2 for the precise definition). Consider
(1) $\mathfrak{N}_{0}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{1,2,3,4,5,6,7,8,9,10,12,13,16,18,25\}.$
Then the coarse moduli space $X_{0}(N)$ of $\mathscr{X}_{0}(N)$ is geometrically rationally connected if and only if it has genus 0, that is if and only if $N\in\mathfrak{N}_{0}$ (c.f. [Maz78]). So we assume that $N\in\mathfrak{N}_{0}$. The $F$-points of these stacks of bounded naive height have been counted by the papers [BS24, Mol23, MV23, AHPP25]. When $N=1$, by the naive height on $\mathscr{X}_{0}(1)\langle F\rangle\cong\mathcal{P}(4,6)\langle F\rangle\to\mathbb{R}_{\geq 0},$ we mean the height function from Definition 3.9. When $F=\mathbb{Q}$, the $\mathbb{Q}$-elliptic curve given by $Y^{2}=X^{3}+eX+f,$ where $e,f\in\mathbb{Z}$ satisfy that for all $p$ prime, one has $p^{4}\nmid e$ or $p^{6}\nmid f$, the naive height is given by $\max(|e^{3}|,|f^{2}|)$. (Note that some sources use a different normalization, e. g. in [AHPP25], it is given by $\max(4|e^{3}|,27|f^{2}|),$ but it has no effect on $a$- and $b$-invariants.) By naive height on $\mathscr{X}_{0}(N)\langle F\rangle$, we mean the composite function $H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathscr{X}_{0}(1)\langle F\rangle\to\mathbb{R}_{>0}$, where $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ forgets the level structure. By Lemma 3.4, one has that $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
Theorem 1.2.
Let $N\in\mathfrak{N}_{0}$ as above and let $M_{\operatorname{naive}}\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ be the element corresponding to the naive height $H$. Assume that $F=\mathbb{Q}$. Then Conjecture 1.1 holds for $\mathscr{X}_{0}(N)$ with the respect to the height $H$.
Theorem 1.3.
For each $N$ and $M_{\operatorname{naive}}$ from Theorem 1.2, one has the following values of $a$- and $b$-invariants:
NN $a(M_{\operatorname{naive}})$ $b(M_{\operatorname{naive}})$
1 5/6 1
2 1/2 1
3 1/3 2
4 1/3 1
5 1/6 5-[F(i):F]
6 1/6 2
7 1/6 2
8 1/6 2
9 1/6 2
10 1/6 1
12 1/6 1
13 1/6 1
16 1/6 1
18 1/6 1
25 1/6 1
Proof of Theorem 1.2 assuming Theorem 1.3.
The theorem follows from Theorem 1.3 and the statements which are listed below:
NN Statement
1 [Bru92, Lemma 4.3]
2 [Gra00, Proposition 1]
3 [BS24, Proposition 4.2]
4 [PS21, Theorem 4.2]
5 [AHPP25, Theorem 1.1]
6 [Mol23, Theorem 1.2.1]
7 [MV23, Theorem 1.2.2]
8 [Mol23, Theorem 1.2.1]
9 [Mol23, Theorem 1.2.1]
10 [Mol23, Theorem 1.2.4]
12 [Mol23, Theorem 1.2.4]
13 [Mol23, Theorem 1.2.4]
16 [Mol23, Theorem 1.2.4]
18 [Mol23, Theorem 1.2.4]
25 [Mol23, Theorem 1.2.4]
Remark 1.4.
Note that when $N=3$, Theorem 1.2 is false without removing a thin set. As it is clear from [BS24, Proposition 4.2], a sufficient thin set is given by
$\{(E,\phi)\in\mathscr{X}_{0}(3)\langle F\rangle|\hskip 2.84544ptj(E)=0\},$
which originated from [PPV20].
1.3. Strategy of proof of Theorem 1.3
To prove Theorem 1.3, we first give a concrete description of $\mathscr{X}_{0}(N)$ as a square root stack over a stacky curve. Recall that given a line bundle $L$ on a stack $\mathcal{X}$, $n$-th root stack of $L$, $\sqrt[n]{L/\mathcal{X}}$, is a stack defined in [Cad07], endowed with a morphism to $\mathcal{X}$ such that the pullback of $L$ is $n$-th power in the Picard group of $\sqrt[n]{L/\mathcal{X}}$.
The square root stacks are $\mu_{2}$-gerbes: they are the trivial $\mu_{2}$-gerbes if and only if $L$ is a square in the Picard group $\operatorname{Pic}(\mathcal{X})$. The square root stacks correspond to $\operatorname{Pic}(X)/2\operatorname{Pic}(X)$. In [AOV08], one describes a notion of rigidification, which essentially kills a subgroup scheme of generic stabilizer group. Hence if $\mathcal{X}$ is a stacky curve and $L$ line bundle, the generic stabilizer is $\mu_{2}$ and the rigidification of $\sqrt[2]{L/\mathcal{X}}$ is just $\mathcal{X}$. We denote the rigidification of $\mathcal{Y}$ by $\mathcal{Y}^{\operatorname{rig}}$.
Similarly, to notion of root stack of a line bundle, there is a notion of $(n_{1},\dots,n_{r})$-th root stack of an $r$-tuple of Cartier divisors $C_{1},\dots,C_{r}$ on an algebraic stack $\mathcal{X}$. When the dimension of $\mathcal{X}$ is equal to $1$, this the stack $\sqrt[n_{1},\dots,n_{r}]{C_{1},\dots,C_{r}/\mathcal{X}}\to X$ can be thought as adding stackiness $\mu_{n_{i}}$ along $C_{i}$. Let $K_{6}$ and $K_{4}$ be the divisors $X^{2}+3Y^{2}=0$ and $X^{2}+Y^{2}=0$, respectively, on $\mathbb{P}^{1}$. Let $E_{0},E_{1728}$ be the divisors corresponding to $X=0$ and $X=1728Y$, respectively.
We have a canonical representable morphism $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)=\mathscr{X}_{0}(1)$ and an induced representable morphism $J_{N}^{\operatorname{rig}}:\mathscr{X}_{0}(N)^{\operatorname{rig}}\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. Let $\Lambda$ be the Hodge line bundle on $\mathscr{X}_{0}(1)$. There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$ such that $\Lambda^{\otimes 2}$ is isomorphic to the pullback of $\Lambda^{\operatorname{rig}}$ under the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. We set $\Lambda_{N}:=J_{N}^{*}\Lambda$ and $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}\Lambda^{\operatorname{rig}}$.
Theorem 1.5.
Assume that $N\in\mathfrak{N}_{0}$. Then there is a canonical isomorphism $\mathscr{X}_{0}(N)\cong\sqrt{\Lambda_{N}^{\operatorname{rig}}/\mathscr{X}_{0}(N)^{\operatorname{rig}}}$ corresponding to the line bundle $\Lambda_{N}$. The following table summarizes the rigidifications of $\mathscr{X}_{0}(N)$, under a suitable isomorphism of $\mathbb{P}^{1}$ and the coarse space $X_{0}(N)$, as root stacks over $\mathbb{P}^{1}$; the closed subschemes $E_{0},E_{1728},K_{4},K_{6}$ of $\mathbb{P}^{1}$ are defined as above.
$\mathscr{X}_{0}(N)$ $\mathscr{X}_{0}(N)^{\operatorname{rig}}$ trivial/non-trivial gerbe
$\mathscr{X}_{0}(1)$ $\sqrt[3,2]{E_{0},E_{1728}/\mathbb{P}^{1}}$ non-trivial
$\mathscr{X}_{0}(2)$ $\sqrt[2]{E_{1728}/\mathbb{P}^{1}}$ non-trivial
$\mathscr{X}_{0}(3)$ $\sqrt[3]{E_{0}/\mathbb{P}^{1}}$ trivial
$\mathscr{X}_{0}(4)$ $\mathbb{P}^{1}$ non-trivial
$\mathscr{X}_{0}(5)$ $\sqrt[2]{K_{4}/\mathbb{P}^{1}}$ non-trivial
$\mathscr{X}_{0}(6)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(7)$ $\sqrt[3]{K_{6}/\mathbb{P}^{1}}$ trivial
$\mathscr{X}_{0}(8)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(9)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(10)$ $\sqrt[2]{K_{4}/\mathbb{P}^{1}}$ non-trivial
$\mathscr{X}_{0}(12)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(13)$ $\sqrt[3,2]{K_{6},K_{4}/\mathbb{P}^{1}}$ non-trivial
$\mathscr{X}_{0}(16)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(18)$ $\mathbb{P}^{1}$ trivial
$\mathscr{X}_{0}(25)$ $\sqrt[2]{K_{4}/\mathbb{P}^{1}}$ non-trivial
The choice of the line bundle $\Lambda_{N}$ induces an identification of automorphism group schemes of points of $\mathscr{X}_{0}(N)$ with $\mu_{m}$ with $m\in\{2,4,6\}$. Let $\mathcal{K}_{6}$ and $\mathcal{K}_{4}$ be the closed substacks of $\mathscr{X}_{0}(N)$ where the automorphism group scheme is of order $6$ and $4$, respectively. Let $I(R)$ be the set of irreducible components of a stack $R$. Using § 2.3.3 (particularly Lemma 2.14), we have the following explicit descriptions of sectors of $\mathscr{X}_{0}(N)$:
Theorem 1.6.
For every $N\in\mathfrak{N}_{0}$, the DM stack $\mathscr{X}_{0}(N)$ has the following sectors:
  • the untwisted sector, that we denote by 0 and which is of age 0;
  • the twisted sector denoted by $1/2$, which is the unique twisted sector of age 0;
  • for $0<\zeta<1$ with $\zeta\in\frac{1}{6}\mathbb{Z}-\frac{1}{2}\mathbb{Z}$ and for $T_{6}\in\mathcal{K}_{6}$, the twisted sector $(T_{6},\zeta)$ which is the equivalence class of the pair $(x,\mu_{6}\xrightarrow{6\zeta}\mu_{6})$ where $x\in T_{6}(\overline{F})$ is a point. The age of $(T_{6},\zeta)$ is $\{4\zeta\}$. Here $\{\cdot\}$ the denotes the fractional part of a real number.
  • for $0<\zeta<1$ with $\zeta\in\frac{1}{4}\mathbb{Z}-\frac{1}{2}\mathbb{Z}$ and for $T_{4}\in\mathcal{K}_{6}$, the twisted sector $(T_{4},\zeta)$ which is the equivalence class of the pair $(x,\mu_{4}\xrightarrow{4\zeta}\mu_{4})$ where $x\in T_{4}(\overline{F})$ is a point. The age of $(T_{4},\zeta)$ is $\{2\zeta\}$.
We identify $\operatorname{NS}(\mathscr{X}_{0}(N))$ with $\mathbb{Q}$ via $\deg:\operatorname{NS}(\mathscr{X}_{0}(N))_{\mathbb{Q}}\xrightarrow{\sim}\mathbb{Q}$. By combining Theorem 1.6 together with additional calculations, we obtain explicit description of $M_{\operatorname{naive}}\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ (see Corollary 3.13) and $K_{\mathscr{X}_{0}(N),\operatorname{orb}}$:
Corollary 1.7.
For every $N\in\mathfrak{N}_{0}$, $K_{\mathscr{X}_{0}(N),\operatorname{orb}}\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ is given by
$((-1+\varepsilon_{2}(N)/4+\varepsilon_{3}(N)/3),(\operatorname{age}(y)-1)_{y\in\pi_{0}^{*}J_{0}\mathscr{X}_{0}(N)})$
where for $u=2,3$, one sets
$\varepsilon_{u}(N):=\begin{cases}0,&\text{ if }u^{2}|N\\ \prod_{p|N}\bigg(1+\bigg(\dfrac{u^{2}-7u+9}{p}\bigg)\bigg),&\text{otherwise.}\end{cases}$
Then Theorem 1.3 reduces down to computations of $a(M_{\operatorname{naive}})$ and $b(M_{\operatorname{naive}})$ based on $M_{\operatorname{naive}},K_{\mathscr{X}_{0}(N),\operatorname{orb}}$ as elements of $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$. The computations of $a(M_{\operatorname{naive}})$ and $b(M_{\operatorname{naive}})$ are divided into two cases: adequate and non-adequate cases in the sense of [DY24, Definition 5.2]. The adequate cases correspond to $N\in\{1,2,3,4,5,6,8,9\}$, where the formula for $a(M_{\operatorname{naive}})$ and $b(M_{\operatorname{naive}})$ are as in [DY24, Conjecture 5.6]; this is the relatively easy case, and is done in Theorem 3.19.
The rest of the cases (when $N\in\{7,10,12,13,16,18,25\}$) correspond to non-adequate cases, where the orbifold pseudoeffective cone $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ plays a key role on computations of $a(M_{\operatorname{naive}})$ and $b(M_{\operatorname{naive}})$. Using (full or partial) description of $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$, explicit values of $a(M_{\operatorname{naive}})$ and $b(M_{\operatorname{naive}})$ are shown in Theorem 4.16 when $N=12,16,18$, Theorem 4.21 when $N=7$, Theorem 4.27 when $N=10,25$, and Theorem 4.30 when $N=13$.
1.4. Organization
In Section 2, we recall cyclotomic stacks, stacky Proj construction, root stacks, and sectors of DM stacks. Section 3 starts with the definition and geometric properties of modular curves $\mathscr{X}_{0}(N)$. After that, the naive height of $\mathscr{X}_{0}(N)$ is studied in Section 3.1, which concludes with the description of the element defining the naive height in $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ (Corollary 3.13). In Section 3.2, we consider the orbifold canonical divisor $K_{\mathscr{X}_{0}(N),\operatorname{orb}}$ and the anticanonical naive height. In Section 3.3, we classify adequate cases and then prove Theorem 1.3 for these cases. Section 4 starts with Section 4.1, where the local computations are discussed to understand stacky curves over $\mathscr{X}_{0}(N)$. Then definitions and properties of the orbifold pseudoeffective cone $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ are described in Section 4.2. In the remaining subsections of Section 4, Theorem 1.3 is proved for the non-adequate cases.
1.5. Notation
Let $F$ be a fixed number field. By $M_{F}$ we denote the set of places of $F$. By $M_{F}^{0}$, respectively by $M_{F}^{\infty}$, we denote the set of finite, respectively, infinite places of $F$. For a finite place $v$ we denote by $q_{v}$ the cardinality of the residue field at $v$. Given a scheme $S$, we will use $(\mu_{n})_{S}$ to denote the finite group scheme of $n$-th roots of unity. If $S=\operatorname{Spec}(K)$, with $K$ a field, we may also write $(\mu_{n})_{K}$ for $(\mu_{n})_{S}$. We may drop $S$ or $K$ from the index, when they are clear from the context. When $K$ a field having all $n$-th roots of unity, we may identify the finite group scheme $(\mu_{n})_{K}$ with the Galois module $\mu_{n}(K)$. By $\zeta_{n}$ we denote a (fixed) generator of $\mu_{n}$. Given a Deligne–Mumford stack $X$, we will denote by $|X|$ the corresponding topological space. Specially, we may denote by $i$ the element $\zeta_{4}$.
1.6. Acknowledgments
The authors would like to thank Mohammad Sadek for many insightful discussions. The authors also would like to thank Santiago Arango-Piñeros, Runxuan Gao, Oana Padurariu, Sun Woo Park, Tim Santens, and Takehiko Yasuda for helpful helpful comments and conversations. The first author has received a funding from the European Union’s Horizon 2023 research and innovation programme under the Maria Skłodowska-Curie grant agreement 101149785. The second author is supported by Korea University grants and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-24535254).
2. Preliminaries
2.1. Cyclotomic stacks and the stacky Proj construction
Let us recall the stacky Proj construction; see [Ols16, Example 10.2.8] or [AHPP25, §2] for details. Given a graded ring $R=\bigoplus_{d\geq 0}R_{d}$, the grading specifies a $\mathbb{G}_{m}$-action on $\operatorname{Spec}R$. This action fixes the point $V(R_{+})$ corresponding to the irrelevant ideal $R_{+}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\bigoplus_{d\geq 0}R_{d}$; when $R_{0}$ is a field or a ring of integers, we abbreviate $V(R_{+})$ as the $\operatorname{Spec}R_{0}$-point 0 in $\operatorname{Spec}R$. Then the stacky Proj of $R$ is the quotient stack
$\operatorname{\mathcal{P}roj}R\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}[(\operatorname{Spec}R-V(R_{+}))/\mathbb{G}_{m}].$
Its coarse moduli space is $\operatorname{Proj}R$, the usual Proj of $R$. On $\operatorname{\mathcal{P}roj}R$, the line bundle $\mathcal{O}_{\operatorname{\mathcal{P}roj}R}(1)$ comes from the associated classifying morphism $\operatorname{\mathcal{P}roj}R\to B\mathbb{G}_{m}$ corresponding to the $\mathbb{G}_{m}$-torsor
$(\operatorname{Spec}R-V(R_{+}))\to\operatorname{\mathcal{P}roj}R.$
If $R$ is a finitely generated graded $F$-algebra with $R_{0}=F$, then $\mathcal{O}_{\operatorname{\mathcal{P}roj}R}(1)$ is ample, i.e. there exists $m>0$ such that $\mathcal{O}_{\operatorname{\mathcal{P}roj}R}(m)$ is isomorphic to the pullback of an ample line bundle on a quasiprojective $F$-variety $\operatorname{Proj}R$.
Example 2.1.
Suppose that a given graded ring $R$ is free and finitely generated in positive degrees, i.e. $R=\mathbb{Z}[x_{0},\dotsc,x_{n}]$ where for every $i$, $x_{i}\in R_{w_{i}}$ for some $w_{i}>0$. Then $\mathcal{P}_{\mathbb{Z}}(w_{0},\dotsc,w_{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\operatorname{\mathcal{P}roj}R$ the weighted projective stack over $\operatorname{Spec}\mathbb{Z}$. Using the definition of $\operatorname{\mathcal{P}roj}$, $\mathcal{P}_{\mathbb{Z}}(w_{0},\dotsc,w_{n})=[(\mathbb{A}^{n+1}_{\mathbb{Z}}-\{0\})/\mathbb{G}_{m}]$, where $\mathbb{G}_{m}$-action on $\mathbb{A}^{n+1}_{\mathbb{Z}}-\{0\}$ is given by $t\cdot(x_{0},\dotsc,x_{n})=(t^{w_{0}}x_{0},\dotsc,t^{w_{n}}x_{n})$. Given a base scheme $S$, $\mathcal{P}_{S}(w_{0},\dotsc,w_{n})$ is the pullback of $\mathcal{P}_{\mathbb{Z}}(w_{0},\dotsc,w_{n})$ on $S$; when there is no ambiguity on $S$, we abbreviate $\mathcal{P}_{S}(w_{0},\dotsc,w_{n})$ as $\mathcal{P}(w_{0},\dotsc,w_{n})$.
Given a $\mathbb{Q}$-stack $\mathcal{X}$, [AHPP25, Lemma 2.7] explains necessary and sufficient conditions for $\mathcal{X}$ to be a stacky Proj, which involves definitions from [AH11, §2]; so let us summarize those definitions over $\operatorname{Spec}F$. A separated $F$-stack $\mathcal{X}$ of locally finitely presentation over $\operatorname{Spec}F$ is called a cyclotomic stack if for every geometric point $\overline{x}$ of $X$, the stabilizer group space $\underline{\operatorname{Aut}}_{\mathcal{X}}(\overline{x})$ (or $\underline{\operatorname{Aut}}(\overline{x})$ for short) is isomorphic to $(\mu_{n})_{\kappa(\overline{x})}$ for some $n\geq 1$. A line bundle $\mathcal{L}$ on a cyclotomic $F$-stack $\mathcal{X}$ is uniformizing if the associated classifying morphism $\mathcal{X}\to B\mathbb{G}_{m}$ is representable; equivalently, for every geometric point $\overline{x}$ of $\mathcal{X}$, $\underline{\operatorname{Aut}}(\overline{x})$-action on $\mathcal{L}_{\overline{x}}$ is faithful (c.f. [AH11, Proposition 2.3.10]). Finally, a uniformizing line bundle $\mathcal{L}$ on a proper cyclotomic $F$-stack $\mathcal{X}$ is polarizing if there exists $m\geq 1$ such that $\mathcal{L}^{m}\cong c^{*}M$, where $c:\mathcal{X}\to X$ is the coarse moduli map and $M$ is an ample line bundle on $X$. A mild generalization of [AHPP25, Lemma 2.7] is as follows:
Lemma 2.2.
Suppose that $\mathcal{X}$ is a proper geometrically connected cyclotomic $F$-stack with a polarizing line bundle $\mathcal{L}$. Then $\mathcal{X}\cong\operatorname{\mathcal{P}roj}R_{\mathcal{L}}$, where $R_{\mathcal{L}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\oplus_{n\geq 0}H^{0}(\mathcal{L}^{n})$ is called the section ring of $\mathcal{L}$.
Proof.
The proof of [AHPP25, Lemma 2.7] follows when $\mathbb{Q}$ is replaced by $F$. ∎
Remark 2.3.
Suppose that $\mathcal{X}$ satisfies the conditions of Lemma 2.2 and the graded $F$-algebra $R_{\mathcal{L}}$ admits a graded resolution by a free graded $F$-algebra $R$ generated in positive degrees. Then the graded resolution $R\twoheadrightarrow R_{\mathcal{L}}$ induces a closed embedding of $\mathcal{X}$, induced by $\mathcal{L}$ into a weighted projective stack $\mathcal{P}(w_{0},\dotsc,w_{n})\cong\operatorname{\mathcal{P}roj}R$ for appropriate $w_{0},\dotsc,w_{n}>0$; in other words, the pullback of $\mathcal{O}_{\mathcal{P}(w_{0},\dotsc,w_{n})}(1)$ on $\mathcal{X}$ is $\mathcal{L}$.
If the generic stabilizer of $\mathcal{X}$ is $\mu_{d}$, then this closed embedding must factor through a coordinate substack $\mathcal{P}(w_{j_{0}},\dotsc,w_{j_{m}})$, where $\{j_{0},\dotsc,j_{m}\}\subset\{0,\dotsc,n\}$ such that $d$ divides $w_{j_{i}}$ for every $0\leq i\leq m$. As a result, $R_{\mathcal{L}}$ is $d\mathbb{N}$-graded.
[…]
Remark 2.4.
[…]
2.2. Root stacks
Here, we summarize the notion of root stacks to describe several classes of cyclotomic $F$-stacks as root stacks. The definitions are from [AGV08, Appendix B] (also independently developed by [Cad07]), but because the definition is local on lisse-étale sites, we will state the construction on stacks.
Given an algebraic $F$-stack $\mathcal{X}$ and a line bundle $\mathcal{L}$, the $n^{\rm th}$ root stack of $\mathcal{L}$ on $\mathcal{X}$ is defined as the fiber product $\sqrt[n]{\mathcal{L}/\mathcal{X}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{X}\times_{B\mathbb{G}_{m}}B\mathbb{G}_{m}$ of $\mathcal{X}\to B\mathbb{G}_{m}$ corresponding to $\mathcal{L}$ and the $n^{\rm th}$-power morphism $B\mathbb{G}_{m}\to B\mathbb{G}_{m}$. Then for every $F$-scheme $T$, any $T$-point of $\sqrt[n]{\mathcal{L}/\mathcal{X}}$ is given by a triple $(x,M,\varphi)$ where $x\colon T\to\mathcal{X}$ is an $F$-morphism, $M$ is a line bundle on $T$, and $\varphi\colon M^{n}\to x^{*}\mathcal{L}$ is an isomorphism. Observe that $\sqrt[n]{\mathcal{L}/\mathcal{X}}$ is a $\mu_{n}$-gerbe over $\mathcal{X}$.
Given an effective Cartier divisor $\mathcal{D}$ on $\mathcal{X}$, pick a global section $s_{\mathcal{D}}$ of $\mathcal{O}_{\mathcal{X}}(\mathcal{D})$ that cuts out $\mathcal{D}$. Then the $n^{\rm th}$ root stack of $\mathcal{D}$ on $\mathcal{X}$ is defined as the fiber product $\sqrt[n]{\mathcal{D}/\mathcal{X}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{X}\times_{[\mathbb{A}^{1}/B\mathbb{G}_{m}]}[\mathbb{A}^{1}/B\mathbb{G}_{m}]$ of $\mathcal{X}\to[\mathbb{A}^{1}/B\mathbb{G}_{m}]$ corresponding to $(\mathcal{L},s_{\mathcal{D}})$ and the $n^{\rm th}$-power morphism $\mathbb{A}^{1}/B\mathbb{G}_{m}]\to[\mathbb{A}^{1}/B\mathbb{G}_{m}]$; the isomorphism class of $\sqrt[n]{\mathcal{L}/\mathcal{X}}$ is independent of the choice of $s_{\mathcal{D}}$. Similarly, given a fixed choice $s_{\mathcal{D}}$, a $T$-object of $\sqrt[n]{\mathcal{D}/\mathcal{X}}$ is instead a quadruple $(x,M,u,\varphi)$ where $u\in H^{0}(M)$ and $\varphi(u^{n})=x^{*}s_{\mathcal{D}}$. Finally, when $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ are mutually disjoint effective Cartier divisors of $\mathcal{X}$, then $\sqrt[n_{1},\dotsc,n_{m}]{\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}/\mathcal{X}}$ means iterative root stack $\sqrt[n_{1}]{\mathcal{D}_{1}/\sqrt[n_{2}]{\mathcal{D}_{2}/\dotsb/\sqrt[n_{m}]{\mathcal{D}_{m}/\mathcal{X}}}}$; note that the isomorphism class of this stack remains the same under permutation of $(n_{1},\mathcal{D}_{1}),\dotsc,(n_{m},\mathcal{D}_{m})$. Because our stacks of interests are DM stacks of dimension one, the mutually disjoint condition of divisors $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ is not restrictive.
Let’s relate root stack construction to rigidification of cyclotomic $F$-stacks:
Lemma 2.5.
Let $\mathcal{X}$ be a cyclotomic $F$-stack with the generic stabilizer $\mu_{n}$, and let $\mathcal{L}$ be a uniformizing line bundle on $\mathcal{X}$. Then $\mathcal{L}$ induces an isomorphism $\mathcal{X}\xrightarrow{\sim}\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$.
Proof.
Observe that the rigidification morphism $\pi\colon\mathcal{X}\to\mathcal{X}^{\operatorname{rig}}$ is a $\mu_{n}$-gerbe by definition. Because $\pi^{*}\mathcal{L}^{\operatorname{rig}}\cong\mathcal{L}^{n}$, this isomorphism induces a morphism $h\colon\mathcal{X}\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$ of $\mu_{n}$-gerbes over $\mathcal{X}^{\operatorname{rig}}$. Then for every smooth $F$-covering $U\to\mathcal{X}^{\operatorname{rig}}$ by an $F$-scheme $U$, the pullback $h_{U}\colon\mathcal{X}\times_{\mathcal{X}^{\operatorname{rig}}}U\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}\times_{\mathcal{X}^{\operatorname{rig}}}U$ is an isomorphism by [Ols16, Lemma 12.2.4], hence $h$ is also an isomorphism by descent. ∎
Corollary 2.6.
Let $\mathcal{X}$ and $\mathcal{Y}$ be geometrically irreducible cyclotomic $F$-stacks with generic stabilizers $\mu_{n}$. Suppose that $f\colon\mathcal{X}\to\mathcal{Y}$ is a representable dominant morphism that induces isomorphism of generic stabilizers. Let $f^{\operatorname{rig}}\colon\mathcal{X}^{\operatorname{rig}}\to\mathcal{Y}^{\operatorname{rig}}$ be the induced morphism of rigidifications. If $\mathcal{L}$ is a uniformizing line bundle on $\mathcal{Y}$, then the following diagram is Cartesian:
$\mathcal{X}$$\mathcal{Y}$$\mathcal{X}^{\operatorname{rig}}$$\mathcal{Y}^{\operatorname{rig}}$$\scriptstyle{f}$$\scriptstyle{\pi_{\mathcal{X}}}$$\scriptstyle{\pi_{\mathcal{Y}}}$$\scriptstyle{f^{\operatorname{rig}}}$
Proof.
Observe that $f^{\operatorname{rig}}$ is dominant and representable. By Lemma 2.5, $\mathcal{Y}\cong\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{Y}^{\operatorname{rig}}}$ and $\pi_{\mathcal{Y}}$ is a $\mu_{n}$-gerbe. As $f^{*}\mathcal{L}$ is a uniformizing line bundle on $\mathcal{X}$, $\pi_{\mathcal{X}}$ is a $\mu_{n}$-gerbe with $\mathcal{X}\cong\sqrt[n]{(f^{*}\mathcal{L})^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$ by Lemma 2.5 again. Note that $(f^{*}\mathcal{L})^{\operatorname{rig}}\cong(f^{\operatorname{rig}})^{*}\mathcal{L}^{\operatorname{rig}}$, which implies that $\mathcal{X}\cong\sqrt[n]{(f^{\operatorname{rig}})^{*}\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$. The formation of root stack commutes with base change, so $\mathcal{X}\cong\mathcal{X}^{\operatorname{rig}}\times_{\mathcal{Y}^{\operatorname{rig}}}\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{Y}^{\operatorname{rig}}}\cong\mathcal{X}^{\operatorname{rig}}\times_{\mathcal{Y}^{\operatorname{rig}}}\mathcal{Y}$. ∎
2.3. Sectors of some Deligne–Mumford curves
2.3.1.
We start by recalling definition of sectors and their ages from [DY24]. Let $X$ be a finite type smooth Deligne–Mumford stack over $X$.
Definition 2.7.
We denote by $\mathcal{J}_{0}X$ its cyclotomic inertia stack, which parametrizes pairs $(x,\mu_{k}\hookrightarrow\operatorname{Aut}(x))$ for some $k\geq 1$. A connected component of $\mathcal{J}_{0}X$ is called a sector of $X$. The connected component corresponding to $k=1$ is called the untwisted sector, while the other components are called twisted sectors.
Definition 2.8.
Let $d\geq 1$ be an integer. Let $\rho$ be a $d$-dimensional representation of a cyclic group $\mu_{n}$, where $n\geq 1$. There exists unique integers $a_{1},\dots,a_{n}\geq 1$ such that $\rho\cong\tau^{a_{1}}\oplus\cdots\tau^{a_{d}}$, with $\tau$ given by the embedding $\mu_{n}\hookrightarrow F^{\times},\zeta_{n}\mapsto\zeta_{n}.$ We define a number
$\operatorname{age}(\rho):=\frac{1}{n}\sum_{i=1}^{d}a_{i}.$
Definition 2.9.
Consider a vector bundle $V$ on $X$. Given an algebraically closed $\overline{F}$-field $K$ an $x\in(\mathcal{J}_{0}X)(K)$ corresponding to $(x^{\prime},\mu_{n}\hookrightarrow\operatorname{Aut}(x))$, we denote by $\rho(x)$ the induced action of $\mu_{n}$ on $V(x)$. We define
$|\mathcal{J}_{0}X|\to\mathbb{Q}_{\geq 0},\hskip 28.45274ptx\mapsto\operatorname{age}(\rho(x)).$
This by [DY24, Lemma 2.22], induces a function
$\operatorname{age}_{V}:\pi_{0}\mathcal{J}_{0}X\to\mathbb{Q}_{\geq 0}.$
If $V=T_{X}$, we write $\operatorname{age}$ for $\operatorname{age}_{V}$.
2.3.2.
The following language is convenient when working with functorial properties. Let us set
[…]
where the transition maps are given by $k|m$ then $\mu_{m}\to\mu_{k}$ is given by $x\mapsto x^{m/k}$. One has that $\widehat{\mu}$ has a structure of a group scheme. By [DY24, Proposition 2.11], one has that the stack $\mathcal{J}_{0}X$ is canonically equivalent with the stack parametrizing the pairs $(x,\widehat{\mu}\to\operatorname{Aut}(x))$ of a point $x$ of $X$ and a homomorphism $\widehat{\mu}\to\operatorname{Aut}(x)$. With this description it is immediate [DY24, Corollary 2.14], that given a morphism $f:Y\to X$ of smooth finite-type Deligne–Mumford stacks, we obtain an induced pointed map $\pi_{0}\mathcal{J}_{0}Y\to\pi_{0}\mathcal{J}_{0}X$.
Definition 2.10.
Let $k$ be an algebraically closed field $\overline{F}$-field. A finite dimensional $k$-representation of $\widehat{\mu}$ is an $k$-homomorphism $\rho:\widehat{\mu}\to\operatorname{GL}_{r}$, for certain $r\geq 1$ called the dimension of representation, which factorizes through the canonical homomorphism $\widehat{\mu}\to\mu_{\ell}$ for certain $\ell\geq 1$. We let $\rho_{\ell}:\mu_{\ell}\to\operatorname{GL}_{r}$ be the induced representation. We define
$\operatorname{age}(\rho):=\operatorname{age}(\rho_{\ell}).$
Lemma 2.11.
Let $\rho$ be an $r$-dimensional $k$-representation. One has that $\operatorname{age}(\rho)$ does not depend on the choice of $\ell$.
Proof.
Choose $\ell$ as in Definition 2.10 minimal possible.. Suppose that $\ell_{0}$ is such that $\rho$ factorizes through $\widehat{\mu}\to\mu_{\ell_{0}}$. Then $\ell|\ell_{0}$. We compare $\operatorname{age}(\rho_{\ell_{0}})$ and $\operatorname{age}(\rho_{\ell})$. Let $\tau_{\ell}:\mu_{\ell}\hookrightarrow\operatorname{GL}_{1}$ and $\tau_{\ell_{0}}:\mu_{\ell}\hookrightarrow\operatorname{GL}_{1}$ be given by the inclusions. Write $\rho_{\ell}$ as
$\rho_{\ell}\cong\bigoplus_{i=1}^{r}\tau_{\ell}^{a_{i}}.$
Then
$\rho_{\ell_{0}}\cong\bigoplus\tau_{\ell_{0}}^{a_{i}\cdot(\ell_{0}/\ell)}.$
We have that
$\operatorname{age}(\rho_{\ell})=\frac{1}{\ell}\sum_{i=1}^{r}a_{i}=\frac{1}{\ell_{0}}\sum_{i=1}^{r}a_{i}\cdot(\ell_{0}/\ell).$
The claim follows. ∎
Lemma 2.12.
Suppose that $f:X\to Y$ is a morphism of smooth finite type $F$-Deligne–Mumford stacks. Let $\mathcal{X}\in\pi_{0}\mathcal{J}_{0}X$ and let $\mathcal{Y}=(\pi_{0}\mathcal{J}_{0}f)(\mathcal{X}).$ Let $V$ be a vector bundle on $Y$. Then
$\operatorname{age}_{f^{*}V}(\mathcal{X})=\operatorname{age}_{V}(\mathcal{Y}).$
Proof.
Let $k$ be an algebraically closed field $\overline{F}$-field such that there exists a $k$-point $x\in\mathcal{X}$. The point $x$ corresponds to a point $x\in X(k)$ and a $k$-homomorphism $\widehat{\mu}\to\operatorname{Aut}(x).$ We have a commutative diagram
$\operatorname{Aut}(x)$$\operatorname{GL}(f^{*}V(x))=\operatorname{GL}(V(f(x))$$\operatorname{Aut}(f(x))$$\operatorname{GL}(V(f(x))$$\scriptstyle{=}$
Let $\rho$ be the induced representation $\rho:\widehat{\mu}\to\operatorname{GL}((f^{*}V)(x))$ from $\operatorname{Aut}(x)\to\operatorname{GL}((f^{*}V)(x))$. Then $\rho$ coincides with the representation $\operatorname{Aut}(x)\to\operatorname{Aut}(f(x))\to\operatorname{GL}(V(f(x)))$. We obtain
$\operatorname{age}_{V}(\mathcal{X})=\operatorname{age}_{f^{*}V}(x)=\operatorname{age}({\rho})=\operatorname{age}_{V}(f(x))=\operatorname{age}_{V}(\mathcal{Y})$
as claimed. ∎
Corollary 2.13.
Suppose we have $f:X\to Y$ as in Lemma 2.12 which is also assumed étale at a neighborhood of a point $x$ of $X$. Let $\mathcal{X}\in\pi_{0}\mathcal{J}_{0}X$ be any sector above $x\in|X|$, and let $\mathcal{Y}=(\pi_{0}\mathcal{J}_{0}f)(\mathcal{X})$. Then
$\operatorname{age}(\mathcal{X})=\operatorname{age}(\mathcal{Y}).$
Proof.
One has that $f$ is étale, then the tangent bundles satisfy $f^{*}T_{Y}=T_{X}$. The corollary follows from Lemma 2.12. ∎
2.3.3. Weighted projective stacks
Let $n\geq 1$ be an integer. Suppose that $a_{1},\dots,a_{n}\geq 1$ are given integers. We consider the weighted projective stack $\mathcal{P}(a_{1},\dots,a_{n}):=[(\mathbb{A}^{n}-\{0\})/\mathbb{G}_{m}]$. For any $F$-scheme $S$ and any morphism $f:S\to\mathcal{P}(a_{1},\dots,a_{n})$, we have a canonical identification $\operatorname{Aut}(f)=\mu_{n(f)}$ for certain $n(f)|\operatorname{lcm}(a_{i}).$ We will fix a bijection bijection
$\delta:\bigcup_{n\geq 1}\operatorname{Hom}(\widehat{\mu},\mu_{n})\xrightarrow{\sim}\mathbb{Q}/\mathbb{Z}.$
Here, for $n|k$, we identify $\theta:\widehat{\mu}\to\mu_{k}$ and $\widehat{\mu}\to\mu_{n}$ which factorizes through $\theta$. If $\widehat{\mu}\to\mu_{n}$ is the canonical surjection, then we define $\delta$ on it to be the class of $1/n$ in $\mathbb{Q}/\mathbb{Z}$. Each homomorphism $\iota:\widehat{\mu}\to\mu_{n}$ is a composite of the canonical homomorphism $\widehat{\mu}\to\mu_{n}$ and a homomorphism $\mu_{n}\to\mu_{n}$ given by $x\mapsto x^{k}$. Then we set $\delta(\iota)=k/n$. The image of $\operatorname{Hom}(\widehat{\mu},\mu_{n})$ is the group $(\frac{1}{n}\mathbb{Z}/\mathbb{Z})$.
Lemma 2.14.
Suppose that $\mathcal{Z}$ is a closed substack of a weighted projective stack. We define a function
$|\mathcal{J}_{0}\mathcal{Z}|\to\mathbb{Q}/\mathbb{Z},\hskip 28.45274pt(x,\phi:\widehat{\mu}\to\operatorname{Aut}(f)=\mu_{n(x)})\mapsto\delta{(\phi)}.$
The function descends to a function $\gamma:\pi_{0}\mathcal{J}_{0}\mathcal{Z}\to\mathbb{Q}/\mathbb{Z}.$
Proof.
Suppose that $S\to\mathcal{J}_{0}\mathcal{Z}$ is a morphism with $S$ connected $F$-scheme corresponding to $(f:S\to\mathcal{Z},\phi:\widehat{\mu}\to\operatorname{Aut}(S)=\mu_{n(f)}).$ Then the value of the function at any geometric point of $S$ is $\delta(\phi)$. The claim follows. ∎
If $z$ is a geometric point of $\mathcal{Z}$ realized as $\widetilde{z}:\operatorname{Spec}(K)\to\mathcal{Z}$ with $K$ an algebraically closed $F$-field, we write $Aut(z~)\operatorname$$\operatorname{Aut}(\widetilde{z})$.
Lemma 2.15.
Denote by $\omega:|\mathcal{J}_{0}\mathcal{Z}|\to|\mathcal{Z}|$ the canonical morphism. Given $z\in|\mathcal{Z}|$, we denote by
$(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}:=\{\mathcal{Y}\in\pi_{0}\mathcal{J}_{0}\mathcal{Z}|\hskip 2.84544pt\exists y\in\mathcal{Y}:\omega(y)=z\}.$
We have a bijection
$\frac{1}{\#\operatorname{Aut}(z)}\mathbb{Z}/\mathbb{Z}\xrightarrow{\sim}(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z},\hskip 28.45274ptx\mapsto[z,\delta^{-1}(x)],$
where $\cdot$] denotes the corresponding connected component.
Proof.
Suppose $z$ is realized as $\widetilde{z}:\operatorname{Spec}(K)\to\mathcal{Z}$ with $K$ an algebraically closed $F$-field. We verify that the map is surjective: if $\mathcal{Y}\in(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}$, let $y\in\mathcal{Y}$ be such that $\omega(y)=z$. By increasing $K$ if needed, we realize $y$ as a $K$-point. The point $y$ then corresponds to a pair $(y_{0},f)$ of a $K$-point of $\mathcal{Z}$, necessarily isomorphic to $\widetilde{z}$, and a homomorphism $f:\widehat{\mu}\to\operatorname{Aut}(y_{0})\cong\operatorname{Aut}(\widetilde{z})=\operatorname{Aut}(z)=\mu_{\#\operatorname{Aut}(z)}$. Now it is $\delta(f)$ that maps to $\mathcal{Y}$. Let us show the injectivity. Suppose $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$ map to the same component. The value of the function $\gamma$ at these components is $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$, respectively. Hence $k_{1}=k_{2}$ and the function is injective. The proof is completed. ∎
3. Modular curves
We recall some properties of the moduli stacks of our interest. Let $\mathscr{Y}_{0}(1)$ be the moduli stack of elliptic curves over $F$. Let $\mathscr{X}_{0}(1)$ be the Deligne–Mumford compactification of $\mathscr{Y}_{0}(1)$. It is well known that
$\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$
and consequently that $\mathscr{X}_{0}(1)^{\operatorname{rig}}\cong\mathcal{P}(2,3).$ The coarse moduli space of $\mathscr{X}_{0}(1)$ is $\mathbb{P}^{1}$. The pullback of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(1)$ is the line bundle $\mathcal{O}_{\mathcal{P}(4,6)}(12)$ on $\mathscr{X}_{0}(1)$, and the Hodge bundle of $\mathscr{X}_{0}(1)$ corresponds to $\mathcal{O}_{\mathcal{P}(4,6)}(1)$, which is polarizing line bundle.
Let $N\geq 1$ be an integer. The following is a special case of [DR73, IV.3.2]:
Definition 3.1.
Let $\mathscr{Y}_{0}(N)$ be the moduli $F$-stack of elliptic curves equipped with cyclic order $N$ subgroup schemes of $E[N]$.
So for every $F$-scheme $S$, every $S$-point of $\mathscr{Y}_{0}(N)$ is of the form $(E\to S,H\subset E[N])$, where $E\to S$ is an elliptic curve defined over $S$ and $H$ is a $S$-flat cyclic subgroup scheme in $E[N]$ of order $N$; so $H$ is étale locally isomorphic to $\mathbb{Z}/N\mathbb{Z}$ as a group scheme. Then any point of $\mathscr{Y}_{0}(N)$ has a non-trivial automorphism group containing $\mu_{2}$ generated by the involution [1][-1] sending $P\in E$ to P-P.
There is a canonical morphism $\mathscr{Y}_{0}(N)\to\mathscr{Y}_{0}(1)$ given by $(E,\phi)\mapsto E$, which forgets the specified cyclic order $N$ subgroup scheme of $E[N]$. This morphism is finite and étale.The open immersion $\mathscr{Y}_{0}(1)\hookrightarrow\mathscr{X}_{0}(1)$ is affine because it is representable and the induced open immersion $Y_{0}(1)\hookrightarrow X_{0}(1)$ of coarse moduli spaces is affine. Thus, $\mathscr{Y}_{0}(N)\to\mathscr{X}_{0}(1)$ is affine as well.
Using the relative normalization of an affine (so representable) morphism $f\colon\mathcal{C}\to\mathcal{S}$ of algebraic stacks as in [Sta26, Tag 0822], we obtain the following definition of $\mathscr{X}_{0}(N)$:
Definition 3.2 ([DR73, IV.3.3]).
Define $\mathscr{X}_{0}(N)$ to be the normalization of $\mathscr{X}_{0}(1)$ with the respect to the affine morphism $\mathscr{Y}_{0}(N)\to\mathscr{Y}_{0}(1)\to\mathscr{X}_{0}(1)$.
Clearly, one can identify $\mathscr{Y}_{0}(N)$ with the preimage of $\mathscr{Y}_{0}(1)$ in $\mathscr{X}_{0}(N)$. We denote by $Y_{0}(N)$ and by $X_{0}(N)$ the coarse moduli spaces of $\mathscr{Y}_{0}(N)$ and by $\mathscr{X}_{0}(N)$, respectively. The curves $Y_{0}(N)$ and $X_{0}(N)$ are well-known to be smooth [DR73, IV.6.7] and irreducible (see discussion in [Yan06, Section 1.1]). Thus $\mathscr{Y}_{0}(N)$ and $\mathscr{X}_{0}(N)$ are irreducible as well.
3.0.1.
Let $\Lambda$ be the Hodge line bundle on $\mathscr{X}_{0}(1)$. There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$, such that its pullback for the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$ is $\Lambda^{2}$. One has a canonical isomorphism $\mathscr{X}_{0}(1)\xrightarrow{\sim}\sqrt[2]{\Lambda^{\operatorname{rig}}/\mathscr{X}_{0}(1)^{\operatorname{rig}}}$ from Lemma 2.5 induced by $\Lambda^{2}\xrightarrow{=}(\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}}).$ Let us write $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ and $\Lambda_{N}:=J_{N}^{*}(\Lambda)$. Let $\eta^{\operatorname{rig}}_{N}:\operatorname{Spec}(\overline{F}(X_{0}(N)))\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ be the generic point of the stacky curve $\mathscr{X}_{0}(N)^{\operatorname{rig}}$.
Lemma 3.3.
The stack $\mathscr{X}_{0}(N)$ is isomorphic to $\operatorname{\mathcal{P}roj}R_{J_{N}^{*}\Lambda}$ and its generic stabilizer is $\mu_{2}$. Furthermore, $\mathscr{X}_{0}(N)\cong\sqrt[2]{\Lambda_{N}^{\operatorname{rig}}/\mathscr{X}_{0}(N)^{\operatorname{rig}}}$.
Proof.
It is clear that generic stabilizer of $\mathscr{X}_{0}(N)$ is $\mu_{2}$. As $\Lambda$ is a polarizing line bundle on $\mathscr{X}_{0}(1)$, we have $\mathscr{X}_{0}(1)\cong\operatorname{\mathcal{P}roj}\Lambda$ by Lemma 2.2. So $\mathscr{X}_{0}(1)\cong\sqrt[2]{\Lambda^{\operatorname{rig}}/\mathscr{X}_{0}(1)^{\operatorname{rig}}}$. Because $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is a finite morphism that induces isomorphism of generic stabilizers, Corollary 2.6 implies that $\mathscr{X}_{0}(N)\cong\mathscr{X}_{0}(N)^{\operatorname{rig}}\times_{\mathscr{X}_{0}(1)^{\operatorname{rig}}}\mathscr{X}_{0}(1)$. Since the formation of root stack commutes with base change, $\mathscr{X}_{0}(N)\cong\sqrt[2]{\Lambda_{N}^{\operatorname{rig}}/\mathscr{X}_{0}(N)^{\operatorname{rig}}}$ by the definition of the polarizing line bundle $\Lambda_{N}$. Therefore, $\mathscr{X}_{0}(N)\cong\operatorname{\mathcal{P}roj}R_{J_{N}^{*}\Lambda}$ by Lemma 2.2. ∎
Lemma 3.4.
One has that $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
Proof.
There exists a non-empty open $U\subset X_{0}(N)\cong\mathbb{P}^{1}$, such that $\mathscr{X}_{0}(N)\cong U\times B\mu_{2}.$ As the fibers of $\mathscr{X}_{0}(N)\to X_{0}(N)$ above points in $U$ are isomorphic to $B\mu_{2}$, they have rational points. The proof follows. ∎
Lemma 3.5.
The degree of $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is equal to the number of cyclic subgroups of order $N$ of $(\mathbb{Z}/N\mathbb{Z})^{2}$. This number is
$\kappa(N):=N\cdot\prod_{p|N}\bigg(1+\frac{1}{p}\bigg).$
Proof.
The first part is clear. To count the number of cyclic subgroups of order $N$, we observe that is precisely $\phi(N)$ times the number of elements of order $N$ in $(\mathbb{Z}/N\mathbb{Z})^{2}$. There are precisely $N\cdot\prod_{p|N}\big(1-\frac{1}{p^{2}}\big)$ elements of order $N$ in this group. The wanted number is thus
$\frac{N\cdot\prod_{p|N}\big(1-\frac{1}{p}^{2}\big)}{N\cdot\prod_{p|N}\big(1-\frac{1}{p}\big)}=\kappa(N).$
By the fact that $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is representable and by [Sil09, Theorem 10.1], the only points of $\mathscr{X}_{0}(N)$ where the automorphism group schemes of points may not be $\mu_{2}$ lie inside of the fibers of $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ where the $j$-value is $1728$ (for the automorphism group scheme subgroup scheme of $\mu_{4}$) and 0 (for the automorphism group scheme subgroup scheme of $\mu_{6}$). The following lemma is a slight restatement of [BV24, Lemma 3.5].
Lemma 3.6.
Let $\phi:R\to R^{\prime}$ be a morphism of DVRs, with ramification index $e$. Let $K$ and $K^{\prime}$ be the the corresponding fields of fractions. Let $n\geq 1$ be an integer. There exists a unique representable morphism
$\sqrt[m]{\operatorname{Spec}(R^{\prime})}\to\sqrt[n]{\operatorname{Spec}(R)}$
extending $\operatorname{Spec}(K^{\prime})\to\operatorname{Spec}(K)$ for some $m\geq 1$. The only $m$ for which the morphism exists is $m=n/\gcd(n,e)$, and in this case its restriction to the closed fiber is the morphism $B\mu_{m}\to B\mu_{n}$, induced by
$\mu_{m}\to\mu_{n},\hskip 28.45274ptx\mapsto x^{e/\gcd(n,e)}.$
Proof.
The existence for $m=n/\gcd(n,e)$ is proven in [BV24, Proof of Lemma 3.5], as well as its description to the closed fiber. The uniqueness is a part of the general result [BV24, Theorem 3.4]. ∎
We end the paragraph by proving the following lemma.
Lemma 3.7.
Suppose that $\widehat{E}\in\mathscr{X}_{0}(N)(\overline{F})$ satisfies that $j(J_{N}(\widehat{E}))\in\{0,1728\}$. The following three properties are equivalent:
  1. (1)
    the map on the coarse moduli spaces $\mathbb{P}^{1}\cong X_{0}(N)\to X_{0}(1)\cong\mathbb{P}^{1}$ is unramified at r(E^)r($\widehat${E}), where $r:\mathscr{X}_{0}(N)\to X_{0}(N)$ is the coarse moduli map;
  2. (2)
    one has $J_{N}$ is étale at $\widehat{E}$ and $\operatorname{Aut}(\widehat{E})=\mu_{6}$ (respectively $\operatorname{Aut}(\widehat{E})=\mu_{4}$) if $j(J_{N}(\widehat{E}))=0$ (respectively $j(J_{N}(\widehat{E}))=1728$);
  3. (3)
[…]
Proof.
[…]
$\sqrt[m_{1}]{\operatorname{Spec}(\widehat{R})}$${\operatorname{Spec}(\widehat{R})}$$\sqrt[m_{2}]{\operatorname{Spec}(R)}$$\operatorname{Spec}(R)$}}
where the left vertical morphism is induced from $J_{N}^{\operatorname{rig}}$ and is representable. By Lemma 3.6 this means that $m_{1}=m_{2}/\gcd(m_{2},e)$ where $e$ is the ramification degree of $R\to\widehat{R}$. By [Sch74, Page 97] one has that $e\in\{1,2,3\}$. As $m_{2}\in\{2,3\}$, this means that that $\overline{J_{N}}$ is unramified at r(E^)r($\widehat${E}) if and only if $m_{1}=m_{2}$, which gives equivalence between 1) and 3). ∎
Lemma 3.8.
There are
$\varepsilon_{3}(N):=\begin{cases}0,&\text{ if }9|N\\ \prod_{p|N}\bigg(1+\bigg(\dfrac{-3}{p}\bigg)\bigg),&\text{otherwise.}\end{cases}$
points in $\mathscr{X}_{0}(N)(\overline{F})$ with automorphism group $\mu_{6}$ and
$\varepsilon_{2}(N):=\begin{cases}0&\text{ if }4|N\\ \prod_{p|N}\bigg(1+\bigg(\dfrac{-1}{p}\bigg)\bigg)&\text{otherwise.}\end{cases}$
points in $\mathscr{X}_{0}(N)(\overline{F})$ with automorphism group $\mu_{4}$. Below we give the table with concrete values:
NN $\varepsilon_{2}(N)$ $\varepsilon_{3}(N)$
1 1 1
2 1 0
3 0 1
4 0 0
5 2 0
6 0 0
7 0 2
8 0 0
9 0 0
10 2 0
12 0 0
13 2 2
16 0 0
18 0 0
25 2 0
Proof.
By Lemma 3.7 we need to count the number of unramified points above $j=0$ and $j=1728$. Now the claim follows from [Sch74, IV, § 7 - 8]. ∎
3.0.2.
We are ready to prove the description of the sectors of $\mathscr{X}_{0}(N)$ from Theorem 1.6. Let
$N\in\{1,2,3,4,5,6,7,8,9,10,12,13,16,18,25\}.$
We have a canonical morphism of stacks $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ which is representable.
Proof of Theorem 1.6.
Let us start by treating the case $N=1$. The choice of the Hodge line bundle $\Lambda$ induces an identification $\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$ with $\Lambda$ corresponding to the line bundle $\mathcal{O}(1)$. We have the untwisted sector $(x,\widehat{\mu}\xrightarrow{y\mapsto 1}\operatorname{Aut}(x)).$ We also have the twisted sector $(x,\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{6}}\mu_{2})$ for any geometric point $x$ with $\operatorname{Aut}(x)=2$. Assume that $\operatorname{Aut}(x)=6,$ hence $\operatorname{Aut}(x)=\mu_{6}$. We have $6$ such sectors by Lemma 2.15, namely the ones containing the points $([0,1],\widehat{\mu}\to\mu_{12}\xrightarrow{x\mapsto x^{2\ell}}\mu_{6})$ for $\ell=0,1,2,3,4,5$. Moreover when $\ell=0,3$, the corresponding sectors are precisely 0 and $1/2$. One treats analogously the sectors corresponding to $\operatorname{Aut}(x)=4.$ Let us calculate the ages. The anticanonical line bundle on $\mathcal{P}(4,6)$ is given by $\mathcal{O}(1)^{\otimes 10}$. The line bundle $\mathcal{O}(1)$ corresponds to the linearization $(x,y,a)\mapsto(t^{4}x,t^{6}y,t^{1}a)$ of the trivial line bundle $(\mathbb{A}^{2}-\{0\})\times\mathbb{A}^{1}\to(\mathbb{A}^{2}-\{0\})$. We deduce that the action of $\mu_{6}$, which is canonically seen as the automorphism group scheme of the point $[0,1]$, on the fiber $\mathcal{O}(1)([0,1])$ is given by $\zeta_{6}\cdot a=\zeta_{6}a.$ In other words, the action corresponds to the canonical inclusion $\mu_{6}\to\overline{F}^{*}=\mathbb{G}_{m}(\overline{F})$. Now we can calculate the age of the sector containing $([0,1],\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{2\ell}}\mu_{6})$. When $\ell=0$, the ages with respect to $\mathcal{O}(1)$ and $\mathcal{O}(1)^{\otimes 10}$ are 0. When $\gcd(\ell,6)=1$, we obtain that the induced inclusion of $\mu_{k}$ for some $k\geq 1$ to the automorphism group scheme of $[0,1]$ is given by $\mu_{6}\to\mu_{6},x\mapsto x^{\ell}$. The age with respect to $\mathcal{O}(1)$ is thus $\ell/6$. It is immediate that the age with respect to $\mathcal{O}(10)$ is $\{10\ell/6\}=\{4\ell/6\}=\{2\ell/3\}$. Assume that $\ell=2$. Then $\widehat{\mu}\to\mu_{6}$ is given by $x\mapsto x^{4}$. The induced inclusion $\mu_{k}\to\mu_{6}$ is given by $\mu_{3}\to\mu_{6},x\mapsto x$. The age with respect to $\mathcal{O}(1)$ is $1/3$, while the age with respect to $\mathcal{O}(10)=\{10/3\}=1/3$. Similarly, in the case $\ell=4$, we get that the induced morphism is $\mu_{3}\to\mu_{6},x\mapsto x^{2}$. The age with respect to $\mathcal{O}(1)$ is $2/3$, while the age with respect to $\mathcal{O}(10)$ is $2/3$. Finally, in the case $\ell=3$, we have that the induced morphism is given by $\mu_{2}\to\mu_{6}$, which has the age $1/2$ with respect to $\mathcal{O}(1)$ and the age 0 with respect to $\mathcal{O}(10)$. The case when the automorphism group is of order $4$ is treated similarly. The proof in the case $N=1$ is completed.
We now treat the general case. The stack $\mathscr{X}_{0}(N)$ has the untwisted sector, that we denote by 0 and which has age 0. Let us choose $x\in\mathscr{X}_{0}(N)(\overline{F})$. Suppose that under $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$, the point $x$ does not map to the points $j=0$ or $j=1728$ of $\mathscr{X}_{0}(1)$. Let $U$ be the complement of the images of $j=0$ and $j=1728$ in $\mathbb{P}^{1}$. The automorphism group of $x$ is isomorphic to $\mu_{2}$. We deduce that $x$ maps to a non-stacky point of $\mathscr{X}_{0}(N)^{\operatorname{rig}}$. Let $s_{2}:\widehat{\mu}\to\mu_{2}$ be the canonical surjection defined over $\operatorname{Spec}(\mathbb{Z})$. The points $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ and $(x,\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ of $\mathcal{J}_{0}\mathscr{X}_{0}(N)(\overline{F})$ map to points of $\mathcal{J}_{0}\mathscr{X}_{0}(1)(\overline{F})$ which lie in distinct connected components. It follows that they lie in distinct connected components of $\mathcal{J}_{0}\mathscr{X}_{0}(N)$. The point $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ lies in the sector 0. We show that $(x_{1},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ and $(x_{2},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$, with $x_{1}$ and $x_{2}$ not mapping to $j=0,1728$, lie in the same component. Indeed, let $\overline{x_{1}}$ and $\overline{x_{2}}$ be the images in $U_{\overline{F}}$ of $x_{1}$ and $x_{2}$. We have that $\mathscr{X}_{0}(N)|_{U}\cong U\times B\mu_{2}$ because $\operatorname{Pic}(U)=\{0\}$. Consider $(U\xrightarrow{(\operatorname{Id}_{U},1)}U\times B\mu_{2}\to\mathscr{X}_{0}(N),((\widehat{\mu})_{U}\xrightarrow{(s_{2})_{U}}(\mu_{2})_{U})$. Restricting it to $\overline{x_{i}}$ gives $(x_{i},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ for $i=1,2$, which means that they are in the same connected component. We call this sector $1/2$. Now it maps to the sector $1/2$ of $\mathscr{X}_{0}(1)$, and Corrolary 2.13 implies that its age is 0.
Consider now a geometric point $x$ above $j=0$ with automorphism group scheme $\mu_{2}$. Let $S$ be the closed substack of $\mathscr{X}_{0}(N)$ with automorphism group scheme larger than $\mu_{2}$. Note that $S$ is finite, i. e. $S\langle\overline{F}\rangle$ is finite. Let $y$ be another finite closed substack of $\mathscr{X}_{0}(N)$ which is geometrically a singleton, distinct from $x$ and distinct from $S$. Set $S^{\prime}:=S\cup\{y\}$. One has that $V:=(\mathscr{X}_{0}(N)-S^{\prime})^{\operatorname{rig}}$ is an open subscheme of $\mathbb{P}^{1}$ and not equal to $\mathbb{P}^{1}$. Hence $\operatorname{Pic}(V)=\{0\}$ and $\mathscr{X}_{0}(N)|_{V}\cong V\times\mathbb{P}^{1}$. Now as above, we have that $(x,(\widehat{\mu})_{\overline{F}}\xrightarrow{s_{\overline{F}}}(\mu)_{\overline{F}})$ is in the sector $1/2$. The same arguments apply if $x$ is a point with automorphism group $\mu_{2}$ lying above $j=1728$.
Consider now a geometric point $x$ above $j=0$, with automorphism group isomorphic to $\mu_{6}$. Clearly, the point $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\operatorname{Aut}(x))$ belongs to the untwisted sector. Consider the point $(x,\widehat{\mu}\xrightarrow{\phi}\operatorname{Aut}(x))$, where $\phi$ is the only homomorphism with the image of size $2$. Let $S_{0}:=S-\{x\}$ and let $S_{0}^{\prime}:=S_{0}\cup\{y\}$ with $y$ as above. Then $(\mathscr{X}_{0}(N)-S_{0}^{\prime})^{\operatorname{rig}}$ is isomorphic to an open substack $\mathcal{U}$ of the stacky curve which is $\mathbb{A}^{1}_{\overline{F}}$ with stackiness $\mu_{3}$ at 0. Let us write $X|_{Y}$ for the fiber product $X\times_{f,Z}Y$ when $f:X\to Z$ is clear from the context. As $\operatorname{Pic}(\mathcal{U})$ is $\mathbb{Z}/3\mathbb{Z}$ [Lop23, Theorem 1.1] in which every element is a square, we obtain that $\mathscr{X}_{0}(N)|_{(\mathscr{X}_{0}(N)-S_{0}^{\prime})^{\operatorname{rig}}}\cong\mathcal{U}\times B\mu_{2}$ so that the point $x$ corresponds to $(0,1)$. The homomorphism $\phi$ becomes $(\widehat{\mu}\xrightarrow{x\mapsto(1,s_{2}(x))}\mu_{3}\times\mu_{2}).$ We remove finitely many non-stacky points of $\mathcal{U}$, if needed, so that $\mathcal{U}$ admits an étale cover by $U^{\prime}$ which is a non-empty open subscheme of $\mathbb{A}^{1}$. We have a morphism $a:U^{\prime}\to\mathcal{U}\to\mathcal{U}\times B\mu_{2}\to\mathscr{X}_{0}(N)$ and $\operatorname{Aut}(a)=\mu_{2}$. Now $(x,\phi)$ is a restriction of $(a,(s_{2})_{U^{\prime}})$ and so is $(x^{\prime},(s_{2})_{\overline{F}})$ for some $x^{\prime}$ with $\operatorname{Aut}(x^{\prime})\cong\mu_{2}$. We deduce that $(x,\phi)$ belongs to the sector $1/2$.
Assume that $x$ is a geometric point of $\mathscr{X}_{0}(N)$ with automorphism group scheme isomorphic to $\mu_{6}$ and that $\widehat{\mu}\to\operatorname{Aut}(x)$ has image of at least $3$ elements. The induced homomorphism $\operatorname{Aut}(x)\to\operatorname{Aut}(J_{N}(x))\cong\mu_{6}$ is an isomorphism. We deduce from Lemma 2.15 that the sectors of $(x,\widehat{\mu}\to\operatorname{Aut}(x))$ are distinct and moreover distinct from the sectors 0 and $1/2$. If $N=3$, there exists precisely one possibility for $x$, so the sectors are described in this case. If $N=7,13$ we have $2$ geometric points of $\mathscr{X}_{0}(N)$ with automorphism group scheme isomorphic to $\mu_{6}$. Moreover, they are defined over $F$ if and only if $\omega\in F$, otherwise they form a connected closed substack of dimension 0. We deduce that in the case $\omega\in F$, we have two sectors of $\mathscr{X}_{0}(N)$ above the sector $j/6$, where $j\in\{1,2,4,5\}$, and if $\omega\not\in F$ we have a single sector above $j/6$. In either case, by Corollary 2.13, the age is equal to the age of the corresponding sector $j/6$ of $\mathscr{X}_{0}(1)$, which is $\{-2j/3\}$ by [DY24, Example 5.9].
[…]
The case $x$ has automorphism group scheme isomorphic to $\mu_{4}$ and $\widehat{\mu}\to\mu_{4}$ is surjective is treated similarly to the case $\widehat{\mu}\to\mu_{6}$ with the image of size at least $3$. The proof is completed. ∎
We may write below for the sector $(\mathcal{K}_{*},t)$ only $t$.
3.1. Naive height
We recall the definition of the naive height $H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathbb{R}_{>0}$. When $N=1$, these heights are called toric heights in [Dar21]. For $v\in M_{F}^{0}$, we define $|\cdot|_{v}$ by $|x|_{v}=q_{v}^{-v(x)}$, where $q_{v}$ is the norm of the maximal ideal $\mathfrak{m}_{v}$ of the ring of integers $\mathcal{O}_{v}$ of $F_{v}$. For $v$ a real place let $|\cdot|_{v}$ be the usual absolute value on $F_{v}\cong\mathbb{R}$. For $v$ a complex place, let $|\cdot|_{v}$ be the square of the absolute value on $F_{v}^{2}\cong\mathbb{C}$. The Hodge line bundle defines an isomorphism $\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$. For $v$ finite, we let
$\mathcal{D}_{v}:=\mathcal{O}_{v}^{2}-\mathfrak{m}_{v}^{4}\times\mathfrak{m}_{v}^{6},$
where $\mathfrak{m}_{v}$ is the maximal ideal of the ring of integers $\mathcal{O}_{v}$ of $F_{v}$. We define a function $r_{v}:F_{v}^{2}-\{0\}\to\mathbb{Z}$, given by
$r_{v}(x_{1},x_{2})=\max\bigg(\bigg\lceil\frac{-v(x_{1})}{4}\bigg\rceil,\bigg\lceil\frac{-v(x_{2})}{6}\bigg\rceil\bigg).$
It follows from [Dar21, Lemma 3.3.4.4 and Lemma 4.4.3.1] that $\pi_{v}^{r_{v}(x_{1},x_{2})}\cdot(x_{1},x_{2})\in\mathcal{D}_{v}$. We set
$f_{v}(x_{1},x_{2})=q_{v}^{-12r_{v}(x_{1},x_{2})}.$
For $v$ infinite we define
$f_{v}:F_{v}^{2}-\{0\}\to\mathbb{R}_{\geq 0},\hskip 28.45274pt(x_{1},x_{2})\mapsto\max(|x_{1}^{3}|_{v},|x_{2}^{2}|_{v}).$
Definition 3.9.
We define a naive height on $\mathscr{X}_{0}(1)$ by
$H:\mathscr{X}_{0}(1)\langle F\rangle\to\mathbb{R}_{>0},\hskip 28.45274pt(x_{1},x_{2})\mapsto\prod_{v\in M_{F}}f_{v}(x_{1},x_{2}).$
We define a naive height on $\mathscr{X}_{0}(N),$ which by abuse of notation is also denoted by $H$:
$H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathbb{R}_{>0},\hskip 28.45274ptx\mapsto H(J_{N}(x)).$
Remark 3.10.
By [Dar21, Example 4.4.3.5], we have when $N=1$ and $F=\mathbb{Q}$, that
$H(x_{1},x_{2})=\max(|x_{1}^{3}|,|x_{2}^{2}|)$
if the coordinates $x_{1},x_{2}$ are in $\mathbb{Z}$ with the condition that for all $p$ prime, one has that $p^{4}\nmid x_{1}\text{ or }p^{6}\nmid x_{2}$.
Let $v$ be a finite place not extending the places $2$ and $3$ and let $K/F_{v}^{\operatorname{un}}$ be the unique extension of degree $12$. The extension is totally ramified and Galois. The Galois group $\operatorname{Gal}(K/F_{v}^{\operatorname{un}})$ canonically identifies with $\mu_{12}$. We deduce a canonical action of $\mu_{12}$ on $\operatorname{Spec}(\mathcal{O}_{K})$. Fix $12$-th root of unity $\zeta_{12}$ and a $12$-th root $\sqrt[12]{\pi_{v}}$ of $\pi_{v}$. Note that the stack $\operatorname{Spec}(\mathcal{O}_{K})/\mu_{12}]$ has one closed point and one open point. The automorphism group of the closed points canonically identifies with $\mu_{12}$ while the open point corresponds to the open immersion $\operatorname{Spec}(F^{\operatorname{un}}_{v})\to[\operatorname{Spec}(\mathcal{O}_{K})/\mu_{12}]$. Let $\psi_{v}:\mathcal{P}(4,6)\langle F_{v}^{\operatorname{un}}\rangle\to\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)$ be the residue map [DY24, Definition 2.17]. Note that the map $\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)\to(\frac{1}{4}\mathbb{Z}\cup\frac{1}{6}\mathbb{Z})\cap[0,1[$ defined by $(\mathcal{K}_{4},t)\mapsto t$, by $(\mathcal{K}_{6},t)\mapsto t$, by $1/2\mapsto 1/2$ and $0\mapsto 0$, is a bijection. We may use this map to denote the sectors.
Lemma 3.11.
Suppose that $x=(x_{1},x_{2})\in\mathscr{X}_{0}(1)\langle F_{v}^{\operatorname{un}}\rangle\cong\mathcal{P}(4,6)\langle F_{v}^{\operatorname{un}}\rangle$. One has that
$\psi_{v}(x_{1},x_{2})=1-\min(v(x_{1})/4,v(x_{2})/6)-r_{v}(x_{1},x_{2}).$
Proof.
One can extend $r_{v}$ to $F_{v}^{\operatorname{un}}$ in the obvious way. We define $\mathcal{D}_{v}^{\operatorname{un}}$ to be $(\mathcal{O}_{v}^{\operatorname{un}})^{2}-\mathfrak{m}_{v}^{\operatorname{un}})^{4}\times(\mathfrak{m}_{v}^{\operatorname{un}})^{6},$ with $\mathcal{O}_{v}$ the ring of the integers in $F_{v}^{\operatorname{un}}$ and $\mathfrak{m}_{v}^{\operatorname{un}}$ its maximal ideal. One has $(x_{1}^{\circ},x_{2}^{\circ}):=\pi_{v}^{r_{v}(x_{1},x_{2})}\cdot(x_{1},x_{2})\in\mathcal{D}^{\operatorname{un}}_{v}$ similarly as above. Let $m=\min({v(x_{1}^{\circ})}/{4},{v(x_{2}^{\circ})}/{6})$. We have that
$m=\min({v(x_{1})}/{4},{v(x_{2})}/{6})+r_{v}(x_{1},x_{2}).$
As $r_{v}(x_{1}^{\circ},x_{2}^{\circ})=0$, it is sufficient to show that
$\psi_{v}(x_{1}^{\circ},x_{2}^{\circ})=1-\min(v(x_{1}^{\circ})/4,v(x_{2}^{\circ})/6)$
for $(x_{1},x_{2})\in\mathcal{D}_{v}^{\operatorname{un}}$. We search for character $\rho:\mu_{12}\to\mathbb{G}_{m}$ corresponding to $\operatorname{Spec}(\mathcal{O}_{K})/\mu_{12}]\to\mathcal{P}(4,6)$ that’s induced by a $\rho$-equivariant morphism $\operatorname{Spec}(\mathcal{O}_{K})\to\mathbb{A}^{2}-\{0\}$ given by $(x_{1},x_{2})$.
Assume that $3v(x_{1})<2v(x_{2})$. The point $(x_{1},x_{2})$ is $K$-isomorphic to the restriction of the point $h:\operatorname{Spec}(\mathcal{O}_{K})\to\mathbb{A}^{2}-\{0\}$ where $h^{*}$ satisfies $X\mapsto(\sqrt[12]{\pi_{v}})^{-48m}x_{1}$, $Y\mapsto(\sqrt[12]{\pi_{v}})^{-72m}x_{2}$. We have a commutative diagram below:
$\sqrt[12]{\pi_{v}}$$((\sqrt[12]{\pi_{v}})^{-48m}x_{1},(\sqrt[12]{\pi_{v}})^{-72m}x_{2})$$\sqrt[12]{\pi_{v}}\zeta_{12}$$((\sqrt[12]{\pi_{v}})^{-48m}\zeta_{12}^{-48m}x_{1},(\sqrt[12]{\pi_{v}})^{-72m}\zeta_{12}^{-72m}x_{2})$$\scriptstyle{h}$$\scriptstyle{h}$
with vertical maps given by $\mu_{12}$ and $\mathbb{G}_{m}$-actions respectively. We deduce $\rho(\zeta_{12})^{4}=\zeta_{12}^{-48m}$ and $\rho(\zeta_{12})^{6}=\zeta_{12}^{-72m}$. Thus $\rho(\zeta_{12})^{2}=\zeta_{12}^{-24m}$ from where we get that $\rho(\zeta_{12})=\zeta_{12}^{-12m}$ or $\rho(\zeta_{12})=\zeta_{12}^{6-12m}$. Note that as the automorphism group of the point $((\sqrt[12]{\pi_{v}})^{-48m}\zeta_{12}^{-48m}x_{1},(\sqrt[12]{\pi_{v}})^{-72m}\zeta_{12}^{-72m}x_{2})$ given by the reduction $(\mathbb{A}^{2}-\{0\})(\mathcal{O}_{K})\to(\mathbb{A}^{2}-\{0\})(\overline{\kappa_{v}})$, where $\overline{\kappa_{v}}$ is the separable closure of the residue field $\kappa_{v}$ at $v$, is $\mu_{6}$. Hence, the induced morphism $\mu_{12}\to\mu_{6}$ is given by $x\mapsto x^{-12m}$. The associated sector is $\{-12m/12\}=1-m$ The case $3v(x_{1})>2v(x_{2})$ is treated similarly. Assume that $3v(x_{1})=2v(x_{2})$. Assume at least on of the inequalities $v(x_{1})<4,v(x_{2})<6$ is valid, we get either $v(x_{1})=v(x_{2})=0$ or $v(x_{1})=2,v(x_{2})=3.$ In the first case the associated sector is the untwisted sector, which is 0 in our notation. In the second case, as above, we get that $\rho(\zeta_{12})=-1$. This implies that the associated sector is $1/2$. The proof is completed. ∎
Lemma 3.12.
The element of $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(1))$ defining the naive height is $(\Lambda,(12-12t)_{t})$ for $t\in\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)$.
Proof.
Consider the morphism $j:\mathscr{X}_{0}(1)\to X_{0}(1)\cong\mathbb{P}^{1}$. The pullback of the line bundle $\mathcal{O}(1)$ is the $12$-th power of the Hodge line bundle $\Lambda^{\otimes 12}$. Endow it with the adelic metric which is the pullback of an adelic metric defining a Weil’s height corresponding to $\mathcal{O}(1)$. Consider the section $j^{*}X$ of $\Lambda^{\otimes 12}$. There exists a finite set of places $S$ containing infinite places and all places above $2$ and $3$, such that for $v\not\in S$, one has whenever $X\neq 0$ that
$||(j^{*}X)(x_{1},x_{2})||_{v}=||X(j(x_{1},x_{2}))||_{v}=|x_{1}^{3}|_{v}\cdot\max(|x_{1}^{3}|_{v},|x_{2}^{2}|_{v})^{-1}=|x_{1}|^{3}_{v}\cdot q_{v}^{12\min(v(x_{1})/4,v(x_{2})/6)}.$
[…]
$(x_{1},x_{2})\mapsto\prod_{v}\frac{{f_{v}(x_{1},x_{2})}}{|x_{1}^{3}|_{v}\cdot\max(|x_{1}^{3}|_{v},|x_{2}^{2}|_{v})}q_{v}^{12-12\psi_{v}(x_{1},x_{2})}=\prod_{v\not\in S}\frac{q_{v}^{-12r_{v}(x_{1},x_{2})+12-12\psi_{v}(x_{1},x_{2})}}{q_{v}^{12\min(v(x_{1})/4,v(x_{2})/6)}}\prod_{v\in S}\frac{f_{v}(x_{1},x_{2})}{\max(|x_{1}^{3}|_{v},|x_{2}^{2}|_{v})^{-1}},$
which is just
$(x_{1},x_{2})\mapsto\prod_{v\in S}\frac{f_{v}(x_{1},x_{2})}{\max(|x_{1}^{3}|_{v},|x_{2}^{2}|_{v}}.$
[…]
Corollary 3.13.
The naive height $H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathbb{R}_{\geq 0}$ corresponds to the element $(J_{N}^{*}\Lambda,(12-12t)_{k\in J_{N}^{-1}(t)})$ where by abuse of notation $J_{N}$ is used for the canonical map $J_{N}:\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(N)\to\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)$. If for some $t\in\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)$ one has that $J_{N}^{-1}(t)=\emptyset$, the corresponding coordinate is understood to be ignored.
3.2. Canonical element
Definition 3.14.
Given a smooth Deligne–Mumford stack $X$ over $F$, we set
$K_{X,\operatorname{orb}}:=(K_{X},(\operatorname{age}(y)-1)_{y\in\pi^{*}_{0}\mathcal{J}_{0}X})$
where $K_{X}$ is the canonical line bundle on $X$ and $\operatorname{age}$ is defined in [DY24, Definition 2.2].
Lemma 3.15.
One has that
$\deg(K_{\mathscr{X}_{0}(N)})=-1+\frac{\varepsilon_{3}(N)}{3}+\frac{\varepsilon_{2}(N)}{4}.$
Proof.
We have a rigidification morphism $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ and $\mathscr{X}_{0}(N)^{\operatorname{rig}}$ has trivial generic stabilizer. Moreover, the coarse moduli morphism $\mathscr{X}_{0}(N)\to X_{0}(N)$ factorizes through $\mathscr{X}_{0}(N)^{\operatorname{rig}}$. Applying [VZB22, Page 58] gives:
$\deg(K_{\mathscr{X}_{0}(N)^{\operatorname{rig}}})$ $=\deg(K_{X_{0}(N)})+\sum_{x\in\mathscr{X}_{0}(N)^{\operatorname{rig}}}\bigg(1-\frac{1}{\#\operatorname{Aut}(x)}\bigg).$
$=-2+\sum_{x\in\mathscr{X}_{0}(N)^{\operatorname{rig}}(\overline{F})}\bigg(1-\frac{1}{\#\operatorname{Aut}(x)}\bigg).$
The rigidification $p:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ is étale by [AOV08, Theorem A1], hence $K_{\mathscr{X}_{0}(N)}=p^{*}K_{\mathscr{X}_{0}(N)^{\operatorname{rig}}}$ and thus $\deg(K_{\mathscr{X}_{0}(N)})=\frac{1}{2}\cdot\deg(K_{\mathscr{X}_{0}(N)^{\operatorname{rig}}})$. We obtain
$\deg(K_{\mathscr{X}_{0}(N)})$ $=\frac{1}{2}\bigg(-2+\sum_{x\in\mathscr{X}_{0}(N)^{\operatorname{rig}}(\overline{F})}\bigg(1-\frac{1}{\#\operatorname{Aut}(x)}\bigg)\bigg)$
$=-1+\sum_{x\in\mathscr{X}_{0}(N)^{\operatorname{rig}}(\overline{F})}\bigg(\frac{1}{2}-\frac{1}{2\#\operatorname{Aut}(x)}\bigg)$
$=-1+\sum_{x\in\mathscr{X}_{0}(N)(\overline{F})}\bigg(\frac{1}{2}-\frac{1}{\#\operatorname{Aut}(x)}\bigg)$
$=-1+\sum_{\begin{subarray}{c}x\in\mathscr{X}_{0}(N)(\overline{F})\\ j(J_{N}(x))=0\end{subarray}}\frac{1}{3}+\sum_{\begin{subarray}{c}x\in\mathscr{X}_{0}(N)(\overline{F})\\ j(J_{N}(x))=1728\end{subarray}}\frac{1}{4}$
$=-1+\frac{\varepsilon_{3}(N)}{3}+\frac{\varepsilon_{2}(N)}{4}.$
Now Lemma 3.7 implies that the number of the points in $\mathscr{X}_{0}(N)(\overline{F})$ with $\operatorname{Aut}(x)=6$ is $\varepsilon_{3}(N)$ while the number of points in $\mathscr{X}_{0}(N)(\overline{F})$ with $\operatorname{Aut}(x)=4$ is $\varepsilon_{2}(N)$. ∎
Lemma 3.16.
The pullback $J_{N}^{*}(\Lambda^{\otimes 12})$ for the canonical morphism $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ has degree
$\deg(J_{N}^{*}(\Lambda^{\otimes 12}))=\frac{\kappa(N)}{2}.$
Proof.
By [DR73, VI.4.4], one has that $\deg(\Lambda)=\frac{1}{24}$. Lemma 3.5 gives
$\deg(J_{N}^{*}(\Lambda^{\otimes 12}))=\kappa(N)\cdot\frac{12}{24}=\frac{\kappa(N)}{2}.$
Definition 3.17.
The anticanonical naive height is the pullback of the naive height, scaled so that the line bundle is the anticanonical line bundle on $\mathscr{X}_{0}(N)$.
Corollary 3.18.
The anticanonical naive height corresponds to the element
$\bigg(1-\frac{\varepsilon_{3}(N)}{3}-\frac{\varepsilon_{2}(N)}{4},\bigg(\frac{24(1-y)\cdot(1-\frac{\varepsilon_{3}(N)}{3}-\frac{\varepsilon_{2}(N)}{4}}{\kappa(N)}\bigg)_{t\in J_{N}^{-1}(y)}\bigg),$
the rational orbifold Néron-Severi space of $\mathscr{X}_{0}(N)$. If for certain $y\in\pi_{0}\mathcal{J}_{0}\mathscr{X}_{0}(1)$ one has that $J_{N}^{-1}(y)$ is empty then the corresponding coordinate is understood to be ignored.
Proof.
This follows immediately from Corollary 3.13 and the previous calculations. ∎
3.3. Adequate cases
Theorem 3.19.
Let $c:\pi_{0}^{*}\mathcal{J}_{0}\mathscr{X}_{0}(N)\to\mathbb{Q}_{\geq 0}$ be the raising function corresponding to the anticanonical naive height. Suppose that for every $y\in\pi_{0}^{*}\mathcal{J}_{0}\mathscr{X}_{0}(N)$, one has that $\operatorname{age}(y)+c(y)\geq 1$. We call this case adequate.
  1. (1)
    The cases $N=1,2,3,4,5,6,8,9$ are adequate.
  2. (2)
    The expected $a$- and $b$-invariants in the adequate cases are given in the table below:
    NN $a$ $b$
    1 5/6 1
    2 1/2 1
    3 1/3 2
    4 1/3 1
    5 1/6 5-[F(i):F]
    6 1/6 2
    8 1/6 2
    9 1/6 2
Proof.
We will use [DY24, Conjecture 5.6] for the adequate cases. The $a$-invariant in this expected to be
$a=\frac{1-\varepsilon_{3}(N)/3-\varepsilon_{2}(N)/4}{\kappa(N)/2},$
while the $b$-invariant is
$b:=\{y\in\pi_{0}^{*}\mathcal{J}_{0}\mathscr{X}_{0}(N)|\hskip 2.84544pt\operatorname{age}(y)+c(y)=1\}+1.$
For $\kappa(N)=N\cdot\prod_{p|N}\big(1+\frac{1}{p}\big),$ we have the following table:
NN $\kappa(N)$
1 1
2 3
3 4
4 6
5 6
6 12
8 12
9 12
The case $N=1$ is adequate by [DY24, Example 5.9]. The corresponding $a$- and $b$-invariants, have been calculated in id. cit. Assume that $N\in\{4,6,8,9\}$. From Lemma 3.8 we see that $\varepsilon_{2}(N)=\varepsilon_{3}(N)=0$. Hence, we have only one twisted sector of age equal to 0. In particular $\kappa(N)\leq 12$ and so $0+24\cdot\frac{1}{2}\cdot\frac{1}{\kappa(N)}\geq 1$ which by Lemma 3.18 implies that $c$ is adequate. In this case we clearly have
$a=\frac{1-\varepsilon_{3}(N)/3-\varepsilon_{2}(N)/4}{\kappa(N)/2}=\frac{2}{\kappa(N)}.$
The $b$-invariant is clearly $1$ if $\kappa(N)<12$ and $2$ if $\kappa(N)=12$. Let us treat the case $N=2$. By Lemma 3.8, we have that $\varepsilon_{2}(2)=1$ and $\varepsilon_{3}(2)=0$. In particular, the only $\overline{F}$-point of $\mathscr{X}_{0}(2)$ with the automorphism group scheme not equal to $\mu_{2}$, has the automorphism group scheme equal to $\mu_{4}$ and is defined over $F$. The $j$-invariant of the corresponding elliptic curve is $1728$. We deduce that $\mathscr{X}_{0}(2)$ has $4$ sectors lying above the sectors $0,1/4,1/2,3/4$ of $\mathscr{X}_{0}(1)$ and having the same ages as them that is, the ages $0,1/2,0,1/2$, respectively. For any twisted sector $t$, we have that
$\operatorname{age}(t)+12\cdot\bigg(\frac{1-1/4}{\kappa(2)/2}\bigg)\cdot(1-J_{2}(t))=\operatorname{age}(t)+6\cdot(1-J_{2}(t))=\begin{cases}3&J_{2}(t)=1/2\\ 5&J_{2}(t)=1/4\\ 2&J_{2}(t)=3/4.\end{cases}$
We obtain that the case $N=2$ is adequate, that $a=(1-1/4)/(\kappa(2)/2)=1/2$ and that $b=1$. Suppose that $N=3$. Then $\varepsilon_{3}(3)=1$, while $\varepsilon_{2}(3)=0$. In particular, the only $\overline{F}$-point of $\mathscr{X}_{0}(3)$ with the automorphism group scheme not equal to $\mu_{2}$, has the automorphism group scheme equal to $\mu_{4}$ and is defined over $F$. The $j$-invariant of the corresponding elliptic curve is 0. We deduce that $\mathscr{X}_{0}(3)$ has $6$ sectors lying above the sectors $0,1/6,1/3,1/2,2/3,5/6$ and having the same ages as them, that is the ages $0,2/3,1/3,0,2/3,1/3$. For any twisted sector $t$, we have that
$\operatorname{age}(t)+12\cdot\bigg(\frac{1-1/3}{\kappa(3)/2}\bigg)\cdot(1-J_{3}(t))=\operatorname{age}(t)+4\cdot(1-J_{3}(t))=\begin{cases}4&J_{3}(t)=1/6\\ 3&J_{3}(t)=1/3\\ 2&J_{3}(t)=1/2\\ 2&J_{3}(t)=2/3\\ 1&J_{3}(t)=5/6.\end{cases}$
We obtain that the case $N=3$ is adequate, that $a=(1-1/3)/(3-\phi(3)/2)=1/3$ and that $b=2$. Assume now $N=5$. By Lemma 3.8, we obtain that $\varepsilon_{2}(5)=2$, while $\varepsilon_{3}(5)=0$. We deduce that the only $\overline{F}$-points of $\mathscr{X}_{0}(5)$ with the automorphism group scheme not equal to $\mu_{2}$ have the automorphism group scheme equal to $\mu_{4}$ and that the $j$-invariant of corresponding elliptic curves is $1728$. Moreover, by [AHPP25, Lemma 2.8], we have that $\mathscr{X}_{0}(N)$ is is isomorphic to the substack of $\mathcal{P}(4,4,2)$ given by $X^{2}+Y^{2}=Z^{4}.$ We deduce that the closed substack given by the points with automorphism group schemes not equal to $\mu_{2},$ that is with the automorphism group scheme isomorphic to $\mu_{4}$, is given by $Z=0$. The two $\overline{F}$-points of this stack are $(i,1,0)$ and (i,1,0)(-i,1,0). We deduce that we have only one closed point of $X_{0}(5)$ lying above the point $j=1728$ of $X_{0}(1)$ if $i\not\in F$ and precisely two points if $i\in F$. Suppose that $i\not\in F$ (respectively, that $i\in F$). We deduce that $\mathscr{X}_{0}(5)$ has $4$ (respectively, $8$) sectors lying above the sectors $0,1/4,1/2,3/4$ of $\mathscr{X}_{0}(1)$ such that above each mentioned sector of $\mathscr{X}_{0}(1)$ there exists precisely $1$ sector (respectively, $2$ sectors, except for the sector $1/2$, above which we have only one sector) of $\mathscr{X}_{0}(5)$. For any twisted sector $t$, we have that
$\operatorname{age}(t)+12\cdot\bigg(\frac{1-1/2}{\kappa(5)/2}\bigg)\cdot(1-J_{5}(t))=\operatorname{age}(t)+2\cdot(1-J_{5}(t))=\begin{cases}5/2,&J_{5}(t)=1/4\\ 1,&J_{5}(t)=1/2\\ 1&J_{5}(t)=3/4.\end{cases}$
Hence $a=1/6$. If $i\not\in F$, we have that $b=3$ and if $i\in F$, then $b=4$. The verification has been completed. ∎
3.3.1.
For the sake of completeness, we express $\mathscr{X}_{0}(2),\mathscr{X}_{0}(3)$ and $\mathscr{X}_{0}(4)$ as root stacks. One has that $\mathscr{X}_{0}(2)^{\operatorname{rig}}\cong\sqrt[2]{E_{1728}}$ where $E_{1728}:X-1728Y=0$, because $\mathscr{X}_{0}(2)^{\operatorname{rig}}$ has only one stacky point which is defined over $F$. In particular, one has that $\mathscr{X}_{0}(2)^{\operatorname{rig}}\cong\mathcal{P}(2,1)$. We deduce from [FMN10, Section 7.3] that $\operatorname{Pic}(\mathscr{X}_{0}(2)^{\operatorname{rig}})=\mathbb{Z}$. Note that $\deg(\operatorname{Pic}(\mathscr{X}_{0}(2)^{\operatorname{rig}}))=\frac{1}{2}\mathbb{Z},$ as every effective Cartier divisor has degree $k/2$ for some $k\geq 0$. Now $(J_{2}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ is not a square, because its degree is $3/2.$ Taking the root with respect to $(J_{2}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ gives:
Corollary 3.20.
One has that $\mathscr{X}_{0}(2)\cong\mathcal{P}(4,2)$ and, in particular, one has that $\mathscr{X}_{0}(2)\to\mathscr{X}_{0}(2)^{\operatorname{rig}}$ is the only (up to isomorphism) non-trivial square root stack over $\mathscr{X}_{0}(2)^{\operatorname{rig}}\cong\mathcal{P}(1,2)$.
Similarly, one has that $\mathscr{X}_{0}(3)^{\operatorname{rig}}\cong\mathcal{P}(1,3)$ hence $\operatorname{Pic}(\mathscr{X}_{0}(3)^{\operatorname{rig}})=\mathbb{Z}$. Note that $\deg(\operatorname{Pic}(\mathscr{X}_{0}(3)^{\operatorname{rig}}))=\frac{1}{3}\mathbb{Z},$ as every effective Cartier divisor has degree $k/3$ for some $k\geq 0$. Now $(J_{3}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ is a square, because its degree is $2/3,$ which is twice the degree of $\mathcal{O}(\mathcal{K}_{3}^{\operatorname{rig}})$.
Corollary 3.21.
One has that $\mathscr{X}_{0}(3)\cong B\mu_{2}\times\mathscr{X}_{0}(3)^{\operatorname{rig}}\cong B\mu_{2}\times\mathcal{P}(1,3)$ and, in particular, that $\mathscr{X}_{0}(3)\to\mathscr{X}_{0}(3)^{\operatorname{rig}}$ is the trivial square root stack.
For $N=4$, we have that $\mathscr{X}_{0}(4)^{\operatorname{rig}}\cong\mathbb{P}^{1}$. Hence one has that $\mathscr{X}_{0}(4)\cong\mathcal{P}(2,2)$ if and only if $(J_{4}^{*})^{\operatorname{rig}}(\Lambda^{\operatorname{rig}})$ is of odd degree. But its degree is $\kappa(4)/6=1$. We obtain:
Corollary 3.22.
One has that $\mathscr{X}_{0}(4)\cong\mathcal{P}(2,2)$ and, in particular, one has that $\mathscr{X}_{0}(4)\to\mathscr{X}_{0}(4)^{\operatorname{rig}}$ is the only (up to isomorphism) non-trivial square root stack over $\mathbb{P}^{1}$.
4. Non-adequate cases
In this section we treat the remaining cases.
4.1. Local situation
We want to treat the other cases. The choice of line bundle $\Lambda_{N}^{\operatorname{rig}}$ on $\mathscr{X}_{0}(N)^{\operatorname{rig}}$ induces an isomorphism of formal neighborhoods of $\overline{F}$-points on $\mathscr{X}_{0}(N)^{\operatorname{rig}}$ with $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$ with $m\in\{1,2,3\}$, where $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$ denotes the $m$-th root stack with respect to the Cartier divisor $m\cdot(T)$. The closed point of $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$ has the automorphism group scheme canonically isomorphic to $\mu_{m}$. Given formal neighborhood $U$ of an $\overline{F}$-point in $\mathscr{X}_{0}(N)^{\operatorname{rig}}$, the choice of line bundle $\Lambda_{N}$ gives an isomorphism $J_{N}^{-1}(U)$ with the root stack $\sqrt{L/\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}}$ with $L$ corresponding to $\Lambda_{N}^{\operatorname{rig}}$.
Let $n\geq 1$. Given a morphism $\operatorname{Spec}(\overline{F}((T)))\to V$ with $V$ some of the $\overline{F}[[T]]$-Deligne–Mumford stacks from below, we find the extension of the morphism $\sqrt[n]{\operatorname{Spec}(\overline{F}[[T]])}\to V$.
Lemma 4.1.
Consider the morphism $\mathbb{A}^{1}\to\mathbb{A}^{1},x\mapsto x^{m}$. Let $x^{m}]:[\mathbb{A}^{1}/\mathbb{G}_{m}]\to[\mathbb{A}^{1}/\mathbb{G}_{m}]$ be the induced morphism from $\mathbb{G}_{m}$-morphism $x\mapsto x^{m}$, where the action of $\mathbb{G}_{m}$ on $\mathbb{A}^{1}$ is $(t,x)\mapsto tx$. The diagram
$\mathcal{P}(1,m)\hskip 11.38092pt$$\mathbb{A}^{1}/\mathbb{G}_{m}]$$\mathbb{P}^{1}$$\mathbb{A}^{1}/\mathbb{G}_{m}]$$\scriptstyle{(\mathcal{O}(1)_{\mathcal{P}(1,m)},x)}$$\scriptstyle{[x^{m}]}$$\scriptstyle{(\mathcal{O}(1),x)}$
is Cartesian.
Proof.
To give a morphism to the fiber product consists to give $(f:T\to\mathbb{P}^{1},L,s,\alpha)$ with $f:T\to\mathbb{P}^{1}$ a morphism, $(L,s)$ a line bundle over $T$ endowed with $s\in\Gamma(T,L)$ and $\alpha:(L^{\otimes m},s^{\otimes m})\xrightarrow{\sim}(f^{*}\mathcal{O}(1),f^{*}(x))$. Let $X\to T$ be the total space of $f^{*}\mathcal{O}(1)$. It is a $\mathbb{G}_{m}$-torsor for the action induced by $\mathbb{G}_{m}\xrightarrow{x\mapsto x^{-1}}\mathbb{G}_{m}$ and the standard action of $\mathbb{G}_{m}$ on $X$. We choose coordinate system $\mathbb{P}^{1}=U\cup V$ with $V:=x\neq 0$ and $U:=y\neq 0$. We have trivializations $(\mathcal{O}(1)|_{V},x|_{V})\xrightarrow{\sim}(V\times\mathbb{A}^{1},1)$ and $(\mathcal{O}(1)|_{U},x|_{U})\xrightarrow{\sim}(U\times\mathbb{A}^{1},1).$ We obtain trivializations $(X|_{f^{-1}(V)},s|_{f^{-1}(V)})\xrightarrow{\sim}(f^{-1}(V)\times\mathbb{A}^{1},1)$ and $(X|_{f^{-1}(U)},s|_{f^{-1}(U)})\xrightarrow{\sim}(f^{-1}(U),1).$ We have two $\mathbb{G}_{m}$-equivariant morphisms $(f^{-1}(V)\times(\mathbb{A}^{1}-\{0\}))\to(\mathbb{A}^{1}-\{0\})\times(\mathbb{A}^{1}-\{0\})$ given by $(r,t)\mapsto(f(r),t)$ and $(f^{-1}(U)\times(\mathbb{A}^{1}-\{0\}))\to(\mathbb{A}^{1}-\{0\})\times(\mathbb{A}^{1}-\{0\})$ given by $(r,t)\mapsto(f(r)^{-1},t).$ They glue to a morphism $X\to\mathbb{A}^{2}-\{0\}$ which is $\mathbb{G}_{m}$-equivariant. Conversely, starting with a morphism $g:T\to\mathcal{P}(1,m)$ with $T$ an $F$-scheme, we take $f:=(\mathcal{P}(1,m)\to\mathbb{P}^{1})\circ g$, $L:=g^{*}(\mathcal{O}(1)_{\mathcal{P}(1,m)})$, $s:=g^{*}x$ and $\alpha$ induced from commutative diagrams. The two processes are inverse one to another. ∎
We obtain a Cartesian diagram:
(2) $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$$\mathcal{P}(1,m)$$\operatorname{Spec}(\overline{F}[[T]])$$\mathbb{P}^{1}$$\scriptstyle{x=0}$$\scriptstyle{x=0}$
The Cartesian diagram (2) identifies the automorphism group of the closed point of $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$ with $\mu_{m}.$
Lemma 4.2.
Let $n\geq 1$. The canonical map
$\operatorname{Pic}\bigg(\sqrt[n]{\operatorname{Spec}(\overline{F}[[T]])}\bigg)\to\operatorname{Pic}(B\mu_{n})=\operatorname{Hom}(\mu_{n},\mathbb{G}_{m})=\mathbb{Z}/n\mathbb{Z}$
[…]
Proof.
Consider the smooth stacky curve $\mathcal{P}(1,n)$. Then $\sqrt[n]{\operatorname{Spec}(\overline{F}[[T]])}$ is the fiber product of the coarse moduli space map and $\operatorname{Spec}(\overline{F}[[T]])\to\mathbb{P}^{1}$ given by the formal neighborhood of 0 in $\mathbb{P}^{1}$. We know by [Lop23, Theorem 1.1] that $\operatorname{Pic}(\mathcal{P}(1,n))\to\mathbb{Z}/n\mathbb{Z}$ is surjective. Clearly this homomorphism factorizes through one in the statement, hence the latter is also surjective. The injectivity follows from [Ols12, Proposition 6.2]. The second claim is immediate. ∎
The diagram (2) induces an equivalence
$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(1,m)}\mathbb{G}_{m}]\xrightarrow{\sim}\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$
where the indices denote the weights of the action. The pullback of the line bundle $\mathcal{O}(1)_{\mathcal{P}(1,m)}$ for
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}\xrightarrow{x=0}\mathcal{P}(1,m)$
induces a line bundle on $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$. The square of the line bundle $\mathcal{O}(1)_{\mathcal{P}(2,m)}$ is isomorphic to the pullback of $\mathcal{O}(1)_{\mathcal{P}(2,3)}$. The pullback of the line bundle $\mathcal{O}(1)_{\mathcal{P}(2,m)}$ for the base change $Z:=\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}\times_{x=0,\mathcal{P}(1,m)}\mathcal{P}(2,m)\to\mathcal{P}(2,m)$ yields an identification
[…]
The following lemmas describe the extension of a morphism $\operatorname{Spec}(\overline{F}((T)))$ to a representable morphism $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$.
Lemma 4.3.
Consider the morphism $\operatorname{Spec}(\overline{F}((T)))\to\operatorname{Spec}(\overline{F}((T)))\times B\mu_{2}$ corresponding to $f\in T\overline{F}[[T]]-\{0\}$ and $g\in H^{1}(\overline{F}((T)),\mu_{2})$. The morphism extends uniquely to a representable morphism of stacks
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}\to\operatorname{Spec}(\overline{F}[[T]])\times B\mu_{2}.$
with $m|2$. Moreover, we have that $m=1$ if and only if $g=0$.
Proof.
The existence and uniqueness follow from [BV24, Theorem 3.1] applied to the commutative diagram:
$\operatorname{Spec}(\overline{F}((T)))$$\operatorname{Spec}(\overline{F}[[T]])\times B\mu_{2}$$\operatorname{Spec}(\overline{F}[[T]])$$\operatorname{Spec}(\overline{F}[[T]]).$}}
We get a representable morphism $\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}\to\operatorname{Spec}(\overline{F}[[T]])\times B\mu_{2}$ which implies that $m|2$. Write
$\operatorname{Spec}(\overline{F}((T)))=[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(0,1)}\mathbb{G}_{m}].$
Write $\operatorname{Spec}(\overline{F}((X)))\times B\mu_{2}=[\operatorname{Spec}(\overline{F}((X))[Y,Y^{-1}])/_{(0,2)}\mathbb{G}_{m}].$ The morphism $(f,g)$ corresponds to $\mathbb{G}_{m}\to\mathbb{G}_{m}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])\to\operatorname{Spec}(\overline{F}((X))[Y,Y^{-1}]),\hskip 28.45274ptX\mapsto f(T),Y\mapsto\widetilde{g}(T)W^{2}$
where $\widetilde{g}(T)=1$ if $g=0$ and $\widetilde{g}(T)=T$ if $g=1$. Endow $\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])$ with the action of $\mathbb{G}_{m}$ given by $t\cdot(T_{0},W_{0})=(t^{1}T_{0},t^{-2}W_{0})$. We have an isomorphism
$\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])\to\operatorname{Spec}(\overline{F}((T))[W,W^{-1}]),\hskip 28.45274ptT\mapsto T_{0}^{2}W_{0}^{1},W\mapsto T_{0}.$
Composing it with $(f,g)$ we get the morphism
$\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])\to\operatorname{Spec}(\overline{F}((X))[Y,Y^{-1}]),\hskip 28.45274ptX\mapsto f(T_{0}^{2}W_{0}^{1}),Y\mapsto\widetilde{g}(T_{0}^{2}W_{0}^{1})T_{0}^{2}.$
Note that $X\mapsto f(T_{0}^{2}W_{0}^{1}),Y\mapsto\widetilde{g}(T_{0}^{2}W_{0}^{1})T_{0}^{2}$ defines a $\mathbb{G}_{m}$-equivariant morphism $\operatorname{Spec}(\overline{F}[[T_{0}]][W_{0},W_{0}^{-1}])\to\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])$. The map on automorphism group schemes at $T_{0}=0$ is given by $\mu_{2}\to\mu_{2},W_{0}\mapsto W_{0}^{1}$ if $g=1$ and $\mu_{2}\to\mu_{2},W_{0}\mapsto 1$ if $g=0$. The claim follows. ∎
We will identify $\mu_{2}\times\mu_{3}$ with $\mu_{6}$ via the isomorphism induced by the canonical inclusions $\mu_{2}\subset\mu_{6}$ and $\mu_{3}\subset\mu_{6}$. We construct an isomorphism $\gamma:\sqrt[3]{\operatorname{Spec}(\overline{F}[[X_{0}]])}\times B\mu_{2}\xrightarrow{\sim}[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(2,6)}\mathbb{G}_{m}].$ Indeed, set $\gamma$ to be the one given by $\mathbb{G}_{m}^{2}\to\mathbb{G}_{m},(t_{0},p_{0})\mapsto t^{2}_{0}p_{0}$-equivariant morphism
$\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1},Z_{0},Z_{0}^{-1}])\to\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}]),\hskip 28.45274ptX\mapsto X_{0}Y_{0}Z_{0},Y\mapsto Y_{0}^{4}Z_{0}^{3}.$
Indeed, let us show that an inverse is provided by $\mathbb{G}_{m}\to\mathbb{G}_{m}^{2},t\mapsto(t^{2},t^{3})$-equivariant morphism
$X_{0}\mapsto X,Y_{0}\mapsto Y,Z_{0}\mapsto Y.$
Indeed, one of the two composites of the morphisms is given by $(X_{0},Y_{0},Z_{0})\mapsto(X_{0}Y_{0}Z_{0},Y_{0}^{4}Z_{0}^{3},Y_{0}^{4}Z_{0}^{3})$ which is isomorphic to the identity via multiplying by $(Y_{0}Z_{0},Y_{0}^{2}Z_{0}).$ The other composite is given by $(X,Y)\mapsto(XY^{2},Y^{7})$ which is isomorphic to the identity via multiplication by $Y.$ Let us find the induced homomorphism on automorphism group schemes of the closed points. Set $X_{0}=0,Y_{0}=1,Z_{0}=1$, the corresponding automorphism group scheme is given by $\mu_{3}\times\mu_{2}$. The image of $(\zeta_{3},\zeta_{2})$ is given by $\zeta_{3}^{4}\zeta_{2}^{2}=\zeta_{3}\zeta_{2},$ hence with the above identification, the induced isomorphism is $\mu_{6}\xrightarrow{=}\mu_{6}$.
Lemma 4.4.
Consider the stack $\sqrt[3]{\operatorname{Spec}(\overline{F}[[T]])}$. Let $f\in T\overline{F}[[T]]-\{0\}$ and $g\in H^{1}(\overline{F}((T)),\mu_{2})$. Let $e=\operatorname{ord}_{T}(f)$. Consider the morphism
$\operatorname{Spec}(\overline{F}((T)))\xrightarrow{(f,g)}\operatorname{Spec}(\overline{F}((T)))\times B\mu_{2}\to\sqrt[3]{\operatorname{Spec}(\overline{F}[[T]])}\times B\mu_{2}$
where the second morphism is the open immersion. The morphism extends uniquely to a representable morphism of stacks
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}\to\sqrt[3]{\operatorname{Spec}(\overline{F}[[T]])}\times B\mu_{2}$
for $m=6/\gcd(6,e)$ if $g=0$ and for $m=6/\gcd(6,3+e)$ if $g=1$. The induced homomorphism is given by $\mu_{m}\to\mu_{6},x\mapsto x^{e}$ if $g=0$ and $x\mapsto x^{e+3}$ if $g=1$.
Proof.
The uniqueness is immediate from the separatedness of the target $\sqrt[3]{\operatorname{Spec}(\overline{F}[[T]])}\times B\mu_{2}$. We will show that there exists an extension $\sqrt[6]{\operatorname{Spec}(\overline{F}[[T]])}\to\sqrt[3]{\operatorname{Spec}(\overline{F}[[T]])}\times B\mu_{2}$ and that the map on the automorphism group scheme is given by $x\mapsto x^{e}$ if $g=0$ and $x\mapsto x^{e+3}$ if $g=1$. We have that
$\operatorname{Spec}(\overline{F}((T)))=[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(0,1)}\mathbb{G}_{m}].$
The morphism $f:\operatorname{Spec}(\overline{F}((T)))\to\operatorname{Spec}(\overline{F}((T)))$ corresponds to $\mathbb{G}_{m}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])\to\operatorname{Spec}(\overline{F}((T))[W,W^{-1}]),\hskip 28.45274ptT\mapsto f(T),W\mapsto W$
Let $\mathbb{G}_{m}$ act on $\operatorname{Spec}(\overline{F}((X_{0}))[Y_{0},Y_{0}^{-1}])$ by $t\cdot(X_{0},Y_{0})=(t^{1}X_{0},t^{3}Y_{0}).$ We have an open immersion
$[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(0,1)}\mathbb{G}_{m}]=\operatorname{Spec}(\overline{F}((T)))$ $\to\sqrt[3]{\operatorname{Spec}(\overline{F}[[X_{0}]])}=[\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1}])/_{(1,3)}\mathbb{G}_{m}],$
$X_{0}\mapsto W,\hskip 2.84544pt$ $Y_{0}\mapsto W^{3}T.$
The morphism $f:\operatorname{Spec}(\overline{F}((T)))\to\operatorname{Spec}(\overline{F}((T)))\to\sqrt[3]{\operatorname{Spec}(\overline{F}[[X_{0}]])}$ corresponds to $\mathbb{G}_{m}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T)))[W,W^{-1}]\to\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1}]),\hskip 28.45274ptX_{0}\mapsto W,Y_{0}\mapsto W^{3}f(T)$
We have an identification
$\sqrt[3]{\operatorname{Spec}(\overline{F}[[X_{0}]]}\times B\mu_{2}=[\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1},Z_{0},Z_{0}^{-1}])/\mathbb{G}_{m}^{2}],\hskip 28.45274pt(t_{0},p_{0})\cdot(X_{0},Y_{0},Z_{0})=(t_{0}^{1}X_{0},t_{0}^{3}Y_{0},p_{0}^{2}Z_{0}).$
The morphism $(f,g):\operatorname{Spec}(\overline{F}((T)))\to[\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1},Z_{0},Z_{0}^{-1}])/\mathbb{G}_{m}^{2}]$ corresponds to the $\mathbb{G}_{m}\to\mathbb{G}_{m}^{2},t\mapsto(t,t)$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])\to\operatorname{Spec}(\overline{F}[[X_{0}]][Y_{0},Y_{0}^{-1},Z_{0},Z_{0}^{-1}]),\hskip 28.45274ptX_{0}\mapsto W,Y_{0}\mapsto W^{3}f(T),Z_{0}\mapsto\widetilde{g}(T)W^{2},$
with $\widetilde{g}(T)=1$ if $g=0$ and $\widetilde{g}(T)=T$ if $g=1$. The composite $\gamma\circ(f,g)$, where $\gamma$ is defined just before the lemma, corresponds to the $\mathbb{G}_{m}\to\mathbb{G}_{m},t\mapsto t^{2}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])\to[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/\mathbb{G}_{m}],\hskip 28.45274ptX\mapsto\widetilde{g}(T)W^{4},Y\mapsto f(T)\widetilde{g}(T)^{3}W^{12}.$
We have an isomorphism
$\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])/\mathbb{G}_{m}]\xrightarrow{\sim}\operatorname{Spec}(\overline{F}((T)))=[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(0,1)}\mathbb{G}_{m}],$
where the action to the left is given by $t\cdot(T_{1},W_{1})=(t^{1}T_{1},t^{6}W_{1})$ induced by the $\mathbb{G}_{m}$-equivariant isomorphism
$\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])\to\operatorname{Spec}(\overline{F}((T))[W,W^{-1}]),\hskip 28.45274ptT\mapsto W_{1}T_{1}^{-6},W\mapsto T_{1}.$
Composing it with $\gamma\circ(f,g)$ yields the morphism $\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])/\mathbb{G}_{m}\to[\operatorname{Spec}(\overline{F}((X))[Y,Y^{-1}])/\mathbb{G}_{m}]$ given by the $\mathbb{G}_{m}\to\mathbb{G}_{m},t\mapsto t^{2}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])\to\operatorname{Spec}(\overline{F}((X))[Y,Y^{-1}]),\hskip 17.07182ptX\mapsto\widetilde{g}(W_{1}T_{1}^{-6})T_{1}^{4},Y\mapsto f(W_{1}T_{1}^{-6})\widetilde{g}(W_{1}T_{1}^{-6})^{3}T_{1}^{12}.$
We write $f(T)=T^{e}f_{0}(T)$ with $f_{0}(T)$ not divisible by $T$. The morphism becomes
$\lambda:X\mapsto\widetilde{g}(W_{1}T_{1}^{-6})T_{1}^{4},\hskip 5.69046ptY\mapsto f_{0}(W_{1}T_{1}^{-6})\cdot(W_{1}T_{1}^{-6})^{e}\widetilde{g}(W_{1}T_{1}^{-6})^{3}T_{1}^{12}=\widetilde{g}(W_{1}T_{1}^{-6})^{3}f_{0}(W_{1}T_{1}^{-6})\cdot W_{1}^{e}T_{1}^{12-6e}.$
If $g=0$, the morphism is $2$-isomorphic to
$X\mapsto T_{1}^{2e},Y\mapsto f_{0}(W_{1}T_{1}^{-6})\cdot W_{1}^{e}$
which extends to a morphism
$\operatorname{Spec}(\overline{F}[[T_{1}]][W_{1},W_{1}^{-1}])/_{(2,6)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(2,6)}\mathbb{G}_{m}].$
$X\mapsto\widetilde{g}(W_{1}T_{1}^{-6})T_{1}^{2e},Y\mapsto\widetilde{g}(W_{1}T_{1}^{-6})f_{0}(W_{1}T_{1}^{-6})\cdot W_{1}^{e}$
with the induced morphism $\mu_{6}\to\mu_{6}$ being given by $W_{1}\mapsto W_{1}^{e}$. Now, we obtain that $m=\gcd(6,6/e)$ and that $\mu_{m}\to\mu_{6}$ is given by $x\mapsto x^{e}$. If $g=1$, that is $\widetilde{g}(W_{1}T_{1}^{-6})=W_{1}T_{1}^{-6}$, the morphism $\lambda$ is given by
$X\mapsto W_{1}T_{1}^{-2},Y\mapsto W_{1}^{3+e}T_{1}^{-18}f_{0}(W_{1}T_{1}^{-6})T_{1}^{12-6e}.$
It is $2$-isomorphic with
$X\mapsto W_{1}T_{1}^{2e},Y\mapsto W_{1}^{3+e}$
which extends to a morphism
$\operatorname{Spec}(\overline{F}[[T_{1}]][W_{1},W_{1}^{-1}])/_{(2,6)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(2,6)}\mathbb{G}_{m}].$
The induced morphism on automorphism group schemes is $\mu_{6}\to\mu_{6},W_{1}\mapsto W_{1}^{3+e}$ if $g=1$. Now we get that $m=6/\gcd(e+3,6)$ and $\mu_{m}\to\mu_{6}$ is given by $x\mapsto x^{e+3}$. The proof is completed. ∎
Thanks to [Lop23, Theorem 1.1], there exists a line bundle on $\mathcal{P}(1,2)$ which maps to a non-trivial line bundle in $\operatorname{Pic}(\sqrt{\operatorname{Spec}(\overline{F}[[X]])})$. Now it is clear that $\mathcal{O}(1)_{\mathcal{P}(1,2)}$ must be such a line bundle. Write $L:=\mathcal{O}(1)|_{\sqrt{\operatorname{Spec}(\overline{F}[[X]])}}$ for the restriction. We have that $\sqrt{\operatorname{Spec}(\overline{F}[[X]])}=[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(1,2)}\mathbb{G}_{m}]$. Consider the stack $\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]$. Let $L_{1}=\mathcal{O}(1)_{[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]}$ be the restriction of $\mathcal{O}(1)_{\mathcal{P}(2,4)}$. The identity $L_{1}^{\otimes 2}=p^{*}L$, where
$p:[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(1,2)}\mathbb{G}_{m}]$
is the canonical morphism, induces an isomorphism
$\sqrt{L/([\operatorname{Spec}(\overline{F}[[X]][Y,Y^{-1}])/_{(1,2)}\mathbb{G}_{m}])}\xrightarrow{\sim}[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}].$
On the automorphism group schemes (as subgroup schemes of $\mathbb{G}_{m}$) at the closed point is given by $\mu_{4}\xrightarrow{=}\mu_{4}$. Furthermore, one has an isomorphism
$\rho:\operatorname{Spec}(\overline{F}((T_{0})))\times B\mu_{2}\xrightarrow{\sim}[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}].$
Indeed, one can identify $\operatorname{Spec}(\overline{F}((T_{0})))\times B\mu_{2}$ with $\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}]/_{(0,2)}\mathbb{G}_{m}]$. Then we have a $\mathbb{G}_{m}$-equivariant isomorphism $\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])\xrightarrow{\sim}\operatorname{Spec}(\overline{F}((T))[W,W^{-1}]$, $(T_{0},W_{0})\mapsto(W_{0},W_{0}^{2}T_{0})$ where the action with weights $(0,2)$ corresponds to the action with the weights $(2,4)$.
Lemma 4.5.
Let $f:=T_{0}^{e}f_{0}(T_{0})$ be such that $e\geq 1$ and $f_{0}$ not divisible by $T$. Let $g\in\mathbb{Z}/2\mathbb{Z}=H^{1}(\overline{F}((T_{0})),\mu_{2})$. The morphism
$\operatorname{Spec}(\overline{F}((T_{0})))\xrightarrow{(f,g)}\operatorname{Spec}(\overline{F}((T_{0})))\times B\mu_{2}\xrightarrow{\rho}\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}\subset[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]$
extends uniquely to a representable morphism
$\lambda:\sqrt[m]{\operatorname{Spec}(\overline{F}[[T_{0}]])}\to[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}],$
for $m=4/\gcd(e,4)$ if $g=0$ and for $m=4/\gcd(e+2,4)$ if $g=1$. The induced homomorphism $\mu_{m}\to\mu_{4}$ is given by $x\mapsto x^{e}$ if $g=0$ and $x\mapsto x^{e+2}$ if $g=1$
Proof.
The uniqueness is clear. With identification $\operatorname{Spec}(\overline{F}((T_{0})))=[\operatorname{Spec}(\overline{F}((T_{0}))[X,X^{-1}])/_{(0,1)}\mathbb{G}_{m}]$, the morphism $(f,g)$ becomes the one induced by the $\mathbb{G}_{m}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T_{0}))[X,X^{-1}])/_{(0,1)}\mathbb{G}_{m}]\rightarrow[\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])/_{(0,2)}\mathbb{G}_{m}],\hskip 19.91684ptT_{0}\mapsto f(T_{0}),W_{0}\mapsto\widetilde{g}(T_{0})X^{2},$
where $\widetilde{g}(T_{0})=1$ if $g=1$ and $\widetilde{g}(T_{0})=T_{0}$ if $g\neq 1$. The morphism $\rho\circ(f,g)$ is given by the $\mathbb{G}_{m}$-equivariant morphism
$\operatorname{Spec}(\overline{F}((T_{0}))[X,X^{-1}])/_{(0,1)}\mathbb{G}_{m}]\rightarrow[{\operatorname{Spec}(\overline{F}((T))[W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}}],\hskip 11.38092ptT\mapsto\widetilde{g}(T_{0})X^{2},W\mapsto\widetilde{g}(T_{0})^{2}X^{4}f(T_{0}).$
Consider the stack $\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])/\mathbb{G}_{m}]$ where the action is $t\cdot(T_{1},W_{1})=(t^{1}T_{1},t^{4}W_{1}).$ This stack is in fact a scheme, and we will provide an isomorphism
$\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])/_{(1,4)}\mathbb{G}_{m}]\xrightarrow{\sim}\operatorname{Spec}\overline{F}((T_{0}))=[\operatorname{Spec}(\overline{F}((T_{0}))[W_{0},W_{0}^{-1}])/_{(0,1)}\mathbb{G}_{m}].$
Indeed, the isomorphism is provided by the $\mathbb{G}_{m}$-equivariant isomorphism
$\operatorname{Spec}(F((T_{1}))[W_{1},W_{1}^{-1}])\xrightarrow{\sim}\operatorname{Spec}(\overline{F}((T_{0}))[X,X^{-1}]),\hskip 28.45274ptT_{0}\mapsto T_{1}^{-4}W_{1}^{1},X\mapsto T_{1}.$
Now $\operatorname{Spec}(\overline{F}((T_{1}))[W_{1},W_{1}^{-1}])/_{(1,4)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}((T))[W,W^{-1}]/_{(2,4)}\mathbb{G}_{m}]$ is given by the $\mathbb{G}_{m}$-equivariant morphism
$T\mapsto\widetilde{g}(T_{1}^{-4}W_{1})T_{1}^{2},W\mapsto\widetilde{g}(T_{1}^{-4}W_{1})^{2}T_{1}^{4}f(T_{1}^{-4}W_{1}).$
We have $f(T_{0})=T_{0}^{e}f_{0}(T_{0})$ with $f_{0}(T_{0})$ not divisible by $T_{0}$. The morphism is $1$-isomorphic to:
$T\mapsto\widetilde{g}(T_{1}^{-4}W_{1}),\hskip 14.22636ptW\mapsto\widetilde{g}(T_{1}^{-4}W_{1})^{2}T_{1}^{-4e}W_{1}^{e}f_{0}(T_{1}^{-4}W_{1}),$
which is $1$-isomorphic with
$T\mapsto\widetilde{g}(T_{1}^{-4}W_{1})\cdot T_{1}^{2e},W\mapsto\widetilde{g}(T_{1}^{-4}W_{1})^{2}W_{1}^{e}f_{0}(T_{1}^{-4}W_{1}).$
Assume that $g=1$, that is $\widetilde{g}=1$. The morphism clearly extends to a morphism
$\operatorname{Spec}([[T_{1}]][W_{1},W_{1}^{-1})/_{(1,4)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]$
and the morphism on the automorphism group schemes at the closed points is given by $\mu_{4}\to\mu_{4},W_{1}\mapsto W_{1}^{e}$. We obtain that $m=4/\gcd(e,2)$ and that $\mu_{m}\to\mu_{4}$ is given by $x\mapsto x$. Assume that $g\neq 1$, that is $\widetilde{g}(T_{0})=T_{0}$. The morphism is $1$-isomorphic to
$T\mapsto W_{1}T_{1}^{2e},W\mapsto W_{1}^{2+e}f_{0}(T_{1}^{-4}W_{1}).$
The morphism clearly extends to a morphism
$\operatorname{Spec}([[T_{1}]][W_{1},W_{1}^{-1})/_{(1,4)}\mathbb{G}_{m}]\to[\operatorname{Spec}(\overline{F}[[T]][W,W^{-1}])/_{(2,4)}\mathbb{G}_{m}]$
and the morphism on the automorphism group schemes at the closed points is given by $\mu_{4}\to\mu_{4},W_{1}\mapsto W_{1}^{2+e}$. We obtain that $m=4/\gcd(e+2,2)$ and that $\mu_{m}\to\mu_{4}$ is given by $x\mapsto x^{-1}$.
The proof is completed.∎
4.1.1. Stacky curves on $\mathscr{X}_{0}(N)$
Let us fix a generic section $\theta_{N}\in\Lambda_{N}$. It induces an isomorphism $\mathcal{O}_{\operatorname{Spec}(\overline{F}((T)))}\xrightarrow{\sim}\Lambda_{N}|_{\eta_{N}^{\operatorname{rig}}}$ and hence an isomorphism $\operatorname{Spec}(\overline{F}(X_{0}(N)))\times B\mu_{2}\xrightarrow{\theta_{N}}\sqrt{\Lambda_{N}^{\operatorname{rig}}|_{\eta^{\operatorname{rig}}_{N}}}$, which, by abuse of notation, we have also denoted by $\theta_{N}$.
Lemma 4.6.
Let $C$ be a smooth projective irreducible $\overline{F}$-curve. Given a pair $(f,g)\in\overline{F}(C)\times H^{1}(\overline{F}(C),\mu_{2})$ with $f$ non-constant, the morphism
$\iota(f,g):\operatorname{Spec}(\overline{F}(C))\xrightarrow{(f,g)}\operatorname{Spec}(\overline{F}(T))\times B\mu_{2}\xrightarrow{\sim}\operatorname{Spec}(\overline{F}(X_{0}(N)))\times B\mu_{2}\xrightarrow{\theta_{N}}\sqrt{\Lambda_{N}^{\operatorname{rig}}|_{\eta^{\operatorname{rig}}_{N}}}\to\sqrt{\Lambda_{N}^{\operatorname{rig}}}\cong\mathscr{X}_{0}(N)$
is affine. The relative normalization $\iota(f,g)^{\nu}:\mathcal{C}\to\mathscr{X}_{0}(N)$ of $\mathscr{X}_{0}(N)$ with respect to $\iota(f,g)$ is representable and $\mathcal{C}$ has $C$ for its coarse moduli space.
Conversely, for any surjective representable $\overline{F}$-morphism $h:\mathcal{C}\to\mathscr{X}_{0}(N)$, with $\mathcal{C}$ a stacky curve having $C$ for its coarse moduli space, there exists a unique pair $(f,g)\in\overline{F}(C)\times H^{1}(\overline{F}(C),\mu_{2})$ with $f$ non-constant, and an isomorphism $\iota(f,g)^{\nu}\implies h$.
Proof.
Let us show the morphism $\iota(f,g)$ is affine. The morphism $g:\operatorname{Spec}(\overline{F}(C))\to B\mu_{2}$ is affine as its base change by a morphism from a scheme to $B\mu_{2}$ is a $\mu_{2}$-torsor over this scheme. The morphism $f$ is affine, hence the product $(f,g)$ is affine. We are left to verify why the inclusion $\sqrt{\Lambda^{\operatorname{rig}}_{\eta_{N}^{\operatorname{rig}}}}\to\sqrt{\Lambda^{\operatorname{rig}}_{N}}$ is affine. This morphism is a base change of the inclusion $\operatorname{Spec}(\overline{F}(X_{0}(N)))\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ so to show it is affine, it will be sufficient to show that the latter morphism is affine. Let $\mathcal{U}\subset\mathscr{X}_{0}(N)^{\operatorname{rig}}$ be the largest open of $\mathscr{X}_{0}(N)$ which is a scheme. Then $\mathcal{U}=\mathscr{X}_{0}(N)^{\operatorname{rig}}\cong\mathbb{P}^{1}$ or $\mathcal{U}$ is affine subvariety of $\mathbb{P}^{1}$. In the first case, the claim follows because the inclusion of the generic point to the projective line is clearly affine (cover the line by two affine open subvarieties) and in the latter case $\operatorname{Spec}(\overline{F}(X_{0}(N)))\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ factorizes through $\mathcal{U}$, hence is affine. Now the first claim follows because relative normalizations are representable morphisms. Let us prove the second claim. Let $\eta_{C}:\operatorname{Spec}(\overline{F}(C))\to\mathcal{C}$ be the generic point. For $(f,g)$ we choose the element given by the identification
$\operatorname{Spec}(\overline{F}(C))\xrightarrow{h|_{\eta_{C}}}\mathscr{X}_{0}(N)|_{\eta^{\operatorname{rig}}_{N}}\cong\sqrt{(\Lambda^{\operatorname{rig}}_{N})_{\eta^{\operatorname{rig}}_{N}}}\xrightarrow{\theta^{-1}_{N}}\operatorname{Spec}(\overline{F}(X_{0}(N)))\times B\mu_{2}\xrightarrow{\sim}\operatorname{Spec}(\overline{F}(T))\times B\mu_{2}.$
Let $\eta_{\mathcal{C}}:\operatorname{Spec}(\overline{F}(C))\to\mathcal{C}$ be the generic point. Now, by the universal property of the relative normalization, there exists an isomorphism $\iota(f,g)^{\nu}\implies h$. ∎
Recall that $\mathcal{K}_{4}$ and $\mathcal{K}_{6}$ are the closed substacks of $\mathscr{X}_{0}(N)$ where the automorphism group is isomorphic to $\mu_{4}$ and $\mu_{6}$, respectively. The two substacks are of dimension 0. One has that $\mathcal{K}_{4}$ (respectively, $\mathcal{K}_{6}$) is connected if and only $i\in F$, (respectively, $\omega\in F$) and in case it is not, then it has precisely two connected components. Write $(\mathcal{K}_{4})_{F(i)}=\mathcal{K}_{4}^{-}\cup\mathcal{K}_{4}^{+}$, (respectively, $(\mathcal{K}_{6})_{\overline{F}(\omega)}=\mathcal{K}_{6}^{-}\cup\mathcal{K}_{6}^{+}$) for the splitting of $\mathcal{K}_{4}$, (respectively, $\mathcal{K}_{6}$) over $F(i)$ (respectively, over $F(\omega)$). Write $K_{4}$ and $K_{6}$ for the closed points of $X_{0}(N)$ corresponding to $\mathcal{K}_{4}$ and $\mathcal{K}_{6}$ for the coarse moduli morphism $\mathscr{X}_{0}(N)\to X_{0}(N)\cong\mathbb{P}^{1}$. Write $(K_{4})_{F(i)}=K_{4}^{-}\cup K_{4}^{+}$ and $(K_{6})_{F(\omega)}=K_{6}^{-}\cup K_{6}^{+}$ with $\mathcal{K}_{i}^{*}$ corresponding to $K_{i}^{*}$ for $(i,*)\in\{4,6\}\times\{+,-\}$. We use the following notation: if $\mathcal{X}$ is a smooth DM stack of dimension $1$ and $x\in\mathcal{X}(\overline{F})$ is a point, we denote by $V(x)$ the formal neighborhood of $x$ in $\mathcal{X}.$ When $\mathcal{X}$ is not a scheme, for our purposes it means the fiber product of the formal neighborhood of the image of $x$ in the coarse moduli space.
Lemma 4.7.
Let $C$ be a smooth projective irreducible $\overline{F}$-curve. Let $(f,g)\in\overline{F}(C)\times H^{1}(\overline{F}(C),\mu_{2})$ with $f$ non-constant and let $\iota(f,g)^{\nu}:\mathcal{C}\to\mathscr{X}_{0}(N)$ be the representable morphism from a smooth stacky curve $\mathcal{C}$ having $C$ for coarse moduli space given by Lemma 4.6. Let $\pi:\mathcal{C}\to C$ be the coarse moduli morphism. For $y\in C(\overline{F})$, denote by $e_{f,y}$ the ramification of $f:C\to\mathbb{P}^{1}\cong X_{0}(N)$ at $y$. The following properties are valid:
  1. (1)
    For any point $x\in C(\overline{F})$ which satisfies that $f(x)\not\in K_{4}\cup K_{6}$, one has that the associated sector of $\iota(f,g)^{\nu}$ at $\pi^{-1}(x)$ is 0 if $g=0\in H^{1}(V(x),\mu_{2})$ and is $1/2$ otherwise;
  2. (2)
    for $\in\{+,-\}$ and for any point $x\in C(\overline{F})$ with $f(x)\in K_{4}^{*}$ , we have that the associated sector of $\iota(f,g)^{\nu}$ at $\pi^{-1}(x)$ is $(\mathcal{K}_{4}^{*},\{(e_{f,x})/4\})$ if $g=0\in H^{1}(V(x),\mu_{2})$ and $(\mathcal{K}_{4}^{*},\{(e_{f,x}+2)/4\})$ otherwise;
  3. (3)
    for $\in\{+,-\}$ and for any point $x\in C(\overline{F})$ with $f(x)\in K_{4}^{*}$, we have that the associated sector of $\iota(f,g)^{\nu}$ at $\pi^{-1}(x)$ is $(\mathcal{K}_{6}^{*},\{(e_{f,x})/6\})$ if $g=0\in H^{1}(V(x),\mu_{2})$ and $(\mathcal{K}_{6}^{*},\{(e_{f,x}+3)/6\})$ otherwise;
Proof.
The choice of the line bundle $J_{N}^{*}(\Lambda)$ on $\mathscr{X}_{0}(N)$ identifies the commutative diagram
$\mathscr{X}_{0}(N)$$\mathscr{X}_{0}(N)^{\operatorname{rig}}$$\mathscr{X}_{0}(1)$$\mathscr{X}_{0}(1)^{\operatorname{rig}}$}}
with the commutative diagram
$\sqrt{(J_{N}^{\operatorname{rig}})^{*}(\mathcal{O}(1)_{\mathscr{P}(2,3)})}$$\sqrt[2,3]{K_{4},K_{6}}$$\mathscr{P}(4,6)$$\mathscr{P}(2,3)$}}
Here the following convention is used: if $K_{6}=\emptyset$, then $\sqrt[2,3]{K_{4},K_{6}}$ is interpreted as $\sqrt[2]{K_{4}}$ and similarly if $K_{4}=\emptyset$. If both are empty simultaneously, we interpret $\sqrt[2,3]{K_{4},K_{6}}$ as $\mathbb{P}^{1}$. By denoting by $r$ the rigidification morphism whatever the origin is, taking the formal neighborhoods yields a commutative diagram
$V(\pi^{-1}(x))$$V({\iota(f,g)^{\nu}(\pi^{-1}(x))})$$V({r(\iota(f,g)^{\nu}(\pi^{-1}(x)))})$$V(J_{N}(\iota(f,g)^{\nu}(\pi^{-1}(x))))$$V(J^{\operatorname{rig}}_{N}(r(\iota(f,g)^{\nu}(\pi^{-1}(x)))))$$\scriptstyle{J_{N}}$$\scriptstyle{J_{N}^{\operatorname{rig}}}$
Assume that $f(x)\not\in K_{4}\cup K_{6}$. The choice of the line bundle $(\mathcal{O}(1)_{\mathcal{P}(4,6)})|_{V(\iota(f,g)^{\nu})}$ yields the diagram
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$$\operatorname{Spec}(\overline{F}[[T]])\times B\mu_{2}$$\operatorname{Spec}(\overline{F}[[T]])$$\operatorname{Spec}(\overline{F}[[T]])\times B\mu_{2}$$\operatorname{Spec}(\overline{F}[[T]])$}}
for some $m|2$. Now observing that the composite of the left horizontal and left vertical morphism is representable and using Lemma 4.3 implies the first claim. Assume that $f(x)\in K_{4}$. The indices in the notation below, when the quotient is taken by $\mathbb{G}_{m}$, represent the weights of the action. The choice of line bundle $(\mathcal{O}(1)_{\mathcal{P}(4,6)})|_{V(\iota(f,g)^{\nu})}$ yields the diagram:
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(2,4)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(1,2)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(2,4)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(1,2)}\mathbb{G}_{m}]$}}
for some $m|4$. The composite of the left horizontal and the left vertical morphism is representable. Using Lemma 4.5 implies that the induced homomorphism on the automorphism group schemes is given by $\mu_{m}\to\mu_{4}$, with $m=1$ if $e\equiv 0\pmod{4}$ and $g=0$ or if $e\equiv 2\pmod{4}$ and $g=1$, with $m=2$ if $e\equiv 2\pmod{4}$ and $g=0$ or $e\equiv 0\pmod{4}$ and $g=1$ and with $m=4$ if $e\equiv 1\pmod{4}$ and $g=0$ or $e\equiv 3\pmod{4}$ and $g=1$. The claim in this case follows. Finally, assume that $f(x)\in K_{6}$. The choice of line bundle $(\mathcal{O}(1)_{\mathcal{P}(4,6)})|_{V(\iota(f,g)^{\nu})}$ yields the diagram:
$\sqrt[m]{\operatorname{Spec}(\overline{F}[[T]])}$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(2,6)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(1,3)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(2,6)}\mathbb{G}_{m}]$$\operatorname{Spec}(\overline{F}[[T]][Y,Y^{-1}])/_{(1,3)}\mathbb{G}_{m}]$}}
for some $m|6$. The composite of the left horizontal and the left vertical morphism is representable. Using Lemma 4.4 and arguing as in the case $m|4$, implies the third claim. ∎
4.2. Orbifold pseudoeffective cone
The orbifold pseudoeffective cone plays the same role in the Batyrev–Manin conjecture for stacks as the pseudoeffective cone plays for varieties. We recall below its definition and properties most of which are established in [DY24]. We then establish some properties which will helps us calculate it for $\mathscr{X}_{0}(N)$ (for some that we need).
Definition 4.8.
[DY24, Definition 8.4] A stacky curve on $\mathscr{X}_{0}(N)_{\overline{F}}$ is a representable morphism $f:\mathcal{C}\to\mathscr{X}_{0}(N)_{\overline{F}}$ with $\mathcal{C}$ a smooth stacky curve. A covering family of stacky curves on $\mathscr{X}_{0}(N)_{\overline{F}}$ is a pair
$(\pi:\widetilde{\mathcal{C}}\to T,\widetilde{f}:\widetilde{\mathcal{C}}\to\mathscr{X}_{0}(N)_{\overline{F}})$
of $\overline{F}$-morphisms of Deligne–Mumford stacks such that:
  • $\pi$ is smooth and surjective,
  • $T$ is an integral scheme of finite type over $\overline{F}$,
  • for each point $t\in T(\overline{F})$, the morphism $\widetilde{f}|_{\pi^{-1}(t)}:\pi^{-1}(t)\to\mathscr{X}_{0}(N)_{\overline{F}}$ is a stacky curve on $\mathscr{X}_{0}(N)_{\overline{F}}$ and
  • $\widetilde{f}$ is dominant.
Definition 4.9.
Let $f:\mathcal{D}\to\mathscr{X}_{0}(N)_{\overline{F}}$ be a morphism with $\mathcal{D}$ a stacky curve, and let $p\in\mathcal{D}(\overline{F})$. We define associated sector of $f$ at $p$ to be the sector of $\mathscr{X}_{0}(N)$ containing the point $(f(p),\widehat{\mu}\to\operatorname{Aut}(p)\to\operatorname{Aut}(f(p)))$. We denote it by $\psi(p)$.
Given $(L,c)\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))_{\mathbb{R}}$ and $f:\mathcal{C}\to\mathscr{X}_{0}(N)_{\overline{F}}$ a stacky curve on $\mathscr{X}_{0}(N)_{\overline{F}}$, we define an intersection number by
$((L,c),f)$ $=\deg(f^{*}L)+\sum_{p\text{ stacky point of }\mathcal{C}}c(\psi(p))$
By [DY24, Lemma 8.6], we have that for a covering family $(\pi:\widetilde{\mathcal{C}}\to T,\widetilde{f}:\widetilde{\mathcal{C}}\to\mathscr{X}_{0}(N)_{\overline{F}})$ and any $(L,c)\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N)_{\overline{F}})$, there exists a non-empty open $T_{0}\subset T$ such that $((L,c),\widetilde{f}|_{\pi^{-1}(t_{0})})$ is constant for $t_{0}\in T_{0}(\overline{F})$. We define $)((L,c),(\pi,\widetilde{f}$)) to be the generic value.
Definition 4.10.
The orbifold pseudo-effective cone of $\mathscr{X}_{0}(N)$ is given by
[…]
Lemma 4.11.
Let $(\pi:\widetilde{\mathcal{C}}\to T,\widetilde{f}:\widetilde{\mathcal{C}}\to\mathscr{X}_{0}(N)_{\overline{F}})$ be a covering family of stacky curves on $\mathscr{X}_{0}(N)_{\overline{F}}$. Then there exists a non-empty open $T_{0}\subset T$ such that 1) for $t\in T_{0}(\overline{F})$ one has that $\widetilde{f}|_{\pi^{-1}(t)}$ surjects onto $\mathscr{X}_{0}(N)_{\overline{F}}$ for $t\in T_{0}(\overline{F})$ or 2) $\widetilde{f}(\pi^{-1}(t))$ is a point for $t\in T_{0}(\overline{F})$. In the latter case, we have that every associated sector of every stacky point of $\widetilde{f}_{\pi^{-1}(t)}$ for $t\in T_{0}(\overline{F})$ is the sector 0 or the sector $1/2$.
Lemma 4.12.
The inequalities $L\geq 0,x_{1/2}\geq 0$ define boundaries of the orbifold pseudoeffective cone $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ For every $(L,c)$ satisfying these inequalities (that is $L\geq 0$ and $c(x_{1/2})\geq 0$) the intersection number with every covering family $(\pi:\widetilde{\mathcal{C}}\to T,\widetilde{f}:\widetilde{\mathcal{C}}\to\mathscr{X}_{0}(N)_{\overline{F}})$ which is of type 2) from Lemma 4.11 is non-negative.
Proof.
Let $x\in\mathscr{X}_{0}(N)_{\overline{F}}(\overline{F})$ be such that $\operatorname{Aut}(x)=\mu_{2}$. By [DY24, Proof of Proposition 9.22], there exists a stacky curve $\mathcal{C}$ and a morphism $g:\mathcal{C}\to(B\mu_{2})_{\operatorname{Spec}(\overline{F})}$ where every associated sector is the non-trivial sector of $B\mu_{2}$. Now for the morphism
$\mathcal{C}\xrightarrow{f}(B\mu_{2})_{\overline{F}}\cong(\operatorname{Spec}(\overline{F})\times_{(\mathscr{X}_{0}(N)_{\overline{F}}\circ\mathscr{X}_{0}(N)^{\operatorname{rig}}_{\overline{F}})\circ x,\mathscr{X}_{0}(N)^{\operatorname{rig}}_{\overline{F}}}\mathscr{X}_{0}(N)_{\overline{F}}\to\mathscr{X}_{0}(N)_{\overline{F}}$
[…]
Remark 4.13.
The orbifold pseudoeffective cone of split toric stacks has been explicitly computed in [DY25a]. When $F=\mathbb{Q}$ and $N\in\{1,2,3,4,6,8,9,12,16,18\}$, the stack $\mathscr{X}_{0}(N)$ is indeed a split toric stack.
4.3. Cases $N=12,16,18$
In this paragraph, we treat the cases $N=12,16,18$.
Theorem 4.14.
For $N\in\{6,8,9,12,16,18\}$, the stacks $\mathscr{X}_{0}(N)$ satisfy that
$\mathscr{X}_{0}(N)\cong\mathbb{P}^{1}\times B\mu_{2}.$
Proof.
For $N$ in the range, we have that $\varepsilon_{2}(N)=\varepsilon_{3}(N)=0$. It follows that the rigidification satisfies $\mathscr{X}_{0}(N)^{\operatorname{rig}}\cong\mathbb{P}^{1}$. We have that $\mathscr{X}_{0}(N)\cong\sqrt[2]{(J_{N}^{\operatorname{rig}})^{*}\Lambda^{\operatorname{rig}}}$. The degree of $\Lambda^{\operatorname{rig}}$ is $\frac{1}{6}$. So we will prove the claim, if we can prove that $(J_{N}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ is a square in the Picard group of $\mathscr{X}_{0}(N)^{\operatorname{rig}}\cong\mathbb{P}^{1}$. Its degree is $\frac{\kappa(N)}{6}$. For $N$ in the range, this is an even number. Hence indeed $(J_{N}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ is a square in $\operatorname{Pic}(\mathscr{X}_{0}(N)^{\operatorname{rig}})$. The claim follows. ∎
Corollary 4.15.
The orbifold pseudoeffective cone of $\mathbb{P}^{1}\times B\mu_{2}$ is given by the inequalities
$\{L\geq 0\},\{x_{1/2}\geq 0\}$
with $L$ denoting the variable corresponding to the numerical class of the line bundle and $x_{1/2}$ the variable corresponding to the only twisted sector.
Proof.
By Lemma 4.12, we have $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))\subset\{L\geq 0,c\geq 0\}$. Consider surjective morphism $f:\mathcal{C}\to\mathscr{X}_{0}(N)\cong\mathbb{P}^{1}\times B\mu_{2}$ with $\mathcal{C}$ a stacky curve. We have that
$((L,c),f)=L+bc(x_{1/2})$
for some $b\geq 0$ an integer. Hence the intersection number is positive. Combining with the last claim of Lemma 4.12 completes the proof. ∎
Theorem 4.16.
The prediction is valid for every of the cases $N\in\{12,16,18\}$ (the other values we have already verified). We have the following values for the $a$- and $b$-invariants:
NN $a$ $b$
12 1/6 1
16 1/6 1
18 1/6 1
Proof.
The naive height corresponds to the element
$(\deg(J_{N}^{*}\Lambda),6)=(\kappa(N)/2,6).$
We have the following table of values for $\kappa(N)$:
NN $\kappa(N)$
12 24
16 24
18 36
The canonical element of $\operatorname{NS}_{\operatorname{orb}}(\mathbb{P}^{1}\times B\mu_{2})$ is given by (1,1)(-1,-1). We have that:
$a=\inf\{t|\hskip 2.84544ptt\cdot(\kappa(N)/2,6)+(-1,-1)\in\mathbb{R}^{2}_{\geq 0}\}=\begin{cases}1/6,&N=12\\ 1/6,&N=16\\ 1/6,&N=18.\end{cases}$
Clearly, the $b$-invariant in each of these cases is $1$. The proof is completed. ∎
4.4. Case $N=7$
We recall the following fact. Let $\omega$ be a primitive third root of unity.
Lemma 4.17.
Let $E_{0}$ be an elliptic curve over $F$ satisfying that $j(E_{0})=0$. One has that
$\operatorname{End}(E_{0})=\begin{cases}\mathbb{Z},&\text{if }\omega\not\in F\\ \mathbb{Z}[\omega],&\text{ otherwise}\end{cases}$
Proof.
One has that $\operatorname{End}((E_{0})_{\overline{F}})=\mathbb{Z}[\omega]$, hence $\mathbb{Z}\subset\operatorname{End}(E_{0})\subset\mathbb{Z}[\omega]$. The absolute Galois group $\operatorname{Gal}(\overline{F}/F)$ acts on $\operatorname{End}((E_{0})_{\overline{F}})$ by $\gamma\cdot\theta=\gamma\circ\theta\circ\gamma^{-1}$. If $\theta$ is represented by $a+b\omega$ with $a,b\in\mathbb{Z}$, the element $\gamma\cdot\theta$ corresponds to
$\gamma\cdot(a+b\omega)=a+b\gamma\cdot\omega=a+b(-1-\omega)=(a-b)-b\omega$
if $\omega\not\in F$ and to
$\gamma\cdot(a+b\omega)=a+b\omega$
if $\omega\in F$. We have that $\operatorname{End}(E_{0})=\operatorname{End}((E_{0})_{\overline{F}})^{\operatorname{Gal}(\overline{F}/F)}$ which is thus $\mathbb{Z}$ iff $\omega\in F$ and $\mathbb{Z}[\omega]$ otherwise. The claim is proven. ∎
Lemma 4.18.
There exist precisely $3-[F(\omega):F]$ closed points of $\mathscr{X}_{0}(7)$ where the automorphism group is $\mu_{6}$ (they are necessarily lying above the point $j=0$ in $\mathscr{X}_{0}(1)$ for the morphism $J_{N}:\mathscr{X}_{0}(7)\to\mathscr{X}_{0}(1)$.)
Proof.
Suppose that $\omega\in F$. Then as $\operatorname{End}(E_{0})=\mathbb{Z}[\omega]$, we deduce that every isogeny of $E_{0}$ is defined over $F$. By Lemma 3.7, we have that the number of isogeny structures is $\varepsilon_{3}(7)=2$. Suppose that $\omega\not\in F$. Then there is no ideal $\mathfrak{a}$ of $\mathbb{Z}$ satisfying that $N_{\mathbb{Q}(\omega)/\mathbb{Q}}(\mathfrak{a})=7$, thus there is no isogeny of degree $7$ defined over $F$. It follows that the number of closed points is $\varepsilon_{3}(7)/2=1$. The proof is completed. ∎
Lemma 4.19.
One has that $\operatorname{Pic}(\mathscr{X}_{0}(7)^{\operatorname{rig}})$ is $\mathbb{Z}$ if $\omega\not\in F$ and $\mathbb{Z}\oplus\mathbb{Z}/3\mathbb{Z}$ otherwise. In particular, if a line bundle $L\in\operatorname{Pic}(\mathscr{X}_{0}(7)^{\operatorname{rig}})$ satisfies that $\deg(L)\in 2\deg(\operatorname{Pic}(\mathscr{X}_{0}(7)^{\operatorname{rig}}))$, then it is a square.
Proof.
We use [Lop23, Theorem 1.1]. Assume first $\omega\not\in F$. Then the quoted theorem gives us the following push-out commutative diagram, where $j_{7}^{*}$ is induced from the coarse moduli space map $j_{7}:\mathscr{X}_{0}(7)^{\operatorname{rig}}\to\mathbb{P}^{1}$:
$\mathbb{Z}^{1}$$\mathbb{Z}^{1}$$\operatorname{Pic}(\mathbb{P}^{1})\cong\mathbb{Z}$$\operatorname{Pic}(\mathscr{X}_{0}(7)^{\operatorname{rig}})$$\scriptstyle{x\mapsto 3x}$$\scriptstyle{x\mapsto 2x}$$\scriptstyle{j_{7}^{*}}$
The pushout is isomorphic to $(\mathbb{Z}\oplus\mathbb{Z})/\{(3m,-2m)|\hskip 2.84544ptm\in\mathbb{Z}\}$ which is $\mathbb{Z}$. Assume that $\omega\in F$. The the quoted theorem gives us the following push-out commutative diagram:
$\mathbb{Z}^{2}$$\mathbb{Z}^{2}$$\operatorname{Pic}(\mathbb{P}^{1})\cong\mathbb{Z}$$\operatorname{Pic}(\mathscr{X}_{0}(3)^{\operatorname{rig}})$$\scriptstyle{(x,y)\mapsto(3x,3y)}$$\scriptstyle{(x,y)\mapsto(x+y)}$$\scriptstyle{j_{7}^{*}}$
We again get $\operatorname{Pic}(\mathscr{X}_{0}(3)^{\operatorname{rig}})$ is $\mathbb{Z}^{3}/\{(3m,3n,-m-n)|\hskip 2.84544ptm,n\in\mathbb{Z}\}$ which is $\mathbb{Z}\oplus(\mathbb{Z}/3\mathbb{Z})$. Let us prove the last statement. Suppose that $\deg(L)=2\deg(L^{\prime})=\deg((L^{\prime})^{\otimes 2})$. We have that $L_{0}:=L\otimes(L^{\prime})^{\otimes-2}$ is of degree 0. There exists an integer $n$ such that $L_{0}^{\otimes n}$ descends to a line bundle on $\mathbb{P}^{1}$. But the degree of the line bundle on $\mathbb{P}^{1}$ is 0, hence it is the trivial line bundle. Hence $L_{0}^{\otimes n}$ is the trivial line bundle, hence $L_{0}$ is a $3$-torsion line bundle. But this means that it is a square. Hence $L=(L^{\prime})^{\otimes 2}\otimes L_{0}$ is a square too. ∎
Theorem 4.20.
Consider the closed point $D:X^{2}+3Y^{2}=0$ of $\mathbb{P}^{1}$. One has that
$\mathscr{X}_{0}(7)\cong\sqrt[3]{D}\times B\mu_{2}.$
Proof.
It follows from Lemma 4.18 that the rigidification is $\mathscr{X}_{0}(7)^{\operatorname{rig}}\cong\sqrt[3]{D}$. The degree of the morphism $J_{7}^{\operatorname{rig}}:\mathscr{X}_{0}(7)^{\operatorname{rig}}\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$ is the same as the degree of $J_{7}:\mathscr{X}_{0}(7)\to\mathscr{X}_{0}(1)$ which is $\kappa(7)=8$. The morphism $J_{7}$ when restricted to $\mathscr{Y}_{0}(7)$ is étale, hence $J_{7}^{\operatorname{rig}}$ is étale when restricted to $J^{\operatorname{rig}}_{7}:\mathscr{Y}_{0}(7)^{\operatorname{rig}}\to\mathscr{Y}_{0}(1)$. The degree of $\mathcal{O}(1)$ on $\mathcal{P}(2,3)$ is $\frac{1}{6}$. The pullback $(J_{7}^{\operatorname{rig}})^{*}(\mathcal{O}(1))$ is of degree $\frac{8}{6}=\frac{4}{3}$. The anticanonical line bundle $K_{\mathscr{X}_{0}(7)^{\operatorname{rig}}}^{-1}$ is of degree $2\cdot\deg(K_{\mathscr{X}_{0}(7)}^{-1})=\frac{2}{3}$. It follows that $(K_{\mathscr{X}_{0}(7)^{\operatorname{rig}}}^{-1})^{\otimes 2}$ and $(J_{7}^{\operatorname{rig}})^{*}(\mathcal{O}(1))$ have the same degree. Hence by Lemma 4.19, we have that $(J_{7}^{\operatorname{rig}})^{*}(\mathcal{O}(1))$ is a square in the Picard group of $\mathscr{X}_{0}(7)^{\operatorname{rig}}$. But $\mathscr{X}_{0}(7)\cong\sqrt[2]{(J_{7}^{\operatorname{rig}})^{*}(\mathcal{O}(1))}$, hence $\mathscr{X}_{0}(7)\cong\sqrt[3]{D}\times B\mu_{2}$. ∎
We will implicitly fix an identification $\mathbb{P}^{1}\cong X_{0}(N)$. Let $(\mathcal{K}_{6})_{\overline{F}}=(\mathcal{K}_{6})^{+}\cup(\mathcal{K}_{6})^{-}$ be the decomposition of $(\mathcal{K}_{6})_{\overline{F}(\omega)}$ into two disjoint points. Denote $(K_{6})^{*}$ the image of $(\mathcal{K}_{6})_{*}$ with $\in\{+,-\}$ for the coarse moduli space map and $K_{6}=(K_{6})^{+}\cup(K_{6})^{-}$. Let $V(x)$ denotes the formal neighborhood of a point $x$ on a curve. Given a smooth projective irreducible $\overline{F}$-curve $C$ and a pair $(f,g)\in(\overline{F}(C)-\overline{F})\times H^{1}(\overline{F}(C),\mu_{2})$ and $(e,i)\in\{0,1,2,3,4,5\}\times\{0,1\}-\{(0,0)\}$, we define $n_{+,e,i}(f,g)$ (respectively, $n_{-,e,i}(f,g)$) to be the number of geometric points $x$ of $C$ which satisfy $f(x)\in(K_{6})^{+}$ (respectively, $f(x)\in(K_{6})^{-}$) and at which the ramification degree of $f$ is congruent to $e$ modulo $6$ and for which $g|_{x}=1\in H^{1}(V_{x},\mu_{2})$ (respectively, $g|_{x}=0\in H^{1}(V_{x},\mu_{2})$) if $i=1$ (respectively, if $i=0$). Specially, we define $n_{-,0,0}(f,g)=n_{+,0,0}(f,g)=0$ and $n_{1/2}(f,g)$ to be the number of geometric points $x$ of $C$ satisfying that $f(x)\not\in K_{6}$ and above which $g$ is ramified. If $k\equiv e\pmod{6}$, by $n_{*,k,j}(f,g)$ we denote $n_{*,e,j}(f,g)$.
It follows from Lemma 4.7, that $\gamma(f,g)\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))^{*}$ is the linear form defined by the representable morphism $\iota(f,g)^{\nu}.$ For a pair $(f,g)$ as before, we denote by $\gamma(f,g)$ the following element of $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}$:
  • Assume that $\omega\not\in F$. We set:
    $\gamma(f,g)(L,x):=\deg(f)\cdot L+\bigg(\sum_{\begin{subarray}{c}1\leq j\leq 5\\ j\neq 3\end{subarray}}x_{(\mathcal{K}_{6},j/6)}\sum_{*\in\{+,-\}}n_{*,j,0}(f,g)+n_{*,3+j,1}(f,g)\bigg)+\\ +x_{1/2}\bigg(n_{1/2}(f,g)+\sum_{*\in\{+,-\}}n_{*,3,0}(f,g)+n_{*,0,1}(f,g)\bigg).$
  • Assume that $\omega\in F$. We set
    $\gamma(f,g)(L,x):=\deg(f)\cdot L+\sum_{*\in\{+,-\}}\sum_{\begin{subarray}{c}1\leq j\leq 5\\ j\neq 3\end{subarray}}x_{((\mathcal{K}_{6})^{*},j/6)}(n_{*,j,0}(f,g)+n_{*,3+j,1}(f,g))+\\ +x_{1/2}\bigg(n_{1/2}(f,g)+\sum_{*\in\{+,-\}}n_{*,3,0}(f,g)+n_{*,0,1}(f,g)\bigg).$
It follows from Lemma 4.7, that $\gamma(f,g)\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))^{*}$ is the linear form defined by the representable morphism $\iota(f,g)^{\nu}.$
Theorem 4.21.
Suppose that $N=7$. The $a$-invariant is given by $1/6.$ The $b$-invariant is given by $b=2.$
Proof.
Using notation $(\mathcal{K}_{6})^{*}$ as before in the case $\omega\in F$ and using it for $\mathcal{K}_{6}$ in the case $\omega\not\in F$, one has that
$K_{\mathscr{X}_{0}(7),\operatorname{orb}}=(-1/3,(-1/3)_{((\mathcal{K}_{6})^{*},1/6)},(-2/3)_{((\mathcal{K}_{6})^{*},1/3)},(-1)_{1/2},(-1/3)_{((\mathcal{K}_{6})^{*},2/3)},(-2/3)_{((\mathcal{K}_{6})^{*},5/6)}).$
The naive height corresponds to
$T:=(4,(10)_{((\mathcal{K}_{6})^{*},1/6)},(8)_{((\mathcal{K}_{6})^{*},1/3)},(6)_{1/2},(4)_{((\mathcal{K}_{6})^{*},2/3)},(2)_{((\mathcal{K}_{6})^{*},5/6)}).$
We have that $x_{1/2}\geq 0$ is an inequality defining cone, so $a\geq 1/6$. Consider the point
$O:$ $=\frac{1}{6}T+K_{\mathscr{X}_{0}(7),\operatorname{orb}}$
$=(2/3,(5/3)_{((\mathcal{K}_{6})^{*},1/6)},(4/3)_{((\mathcal{K}_{6})^{*},1/3)},(1)_{1/2},(2/3)_{((\mathcal{K}_{6})^{*},2/3)},(1/3)_{((\mathcal{K}_{6})^{*},5/6)})+$
$\quad\quad+(-1/3,(-1/3)_{((\mathcal{K}_{6})^{*},1/6)},(-2/3)_{((\mathcal{K}_{6})^{*},1/3)},(-1)_{1/2},(-1/3)_{((\mathcal{K}_{6})^{*},2/3)},(-2/3)_{((\mathcal{K}_{6})^{*},5/6)})$
$=(1/3,(4/3)_{((\mathcal{K}_{6})^{*},1/6)},(2/3)_{((\mathcal{K}_{6})^{*},1/3)},(0)_{1/2},(1/3)_{((\mathcal{K}_{6})^{*},2/3)},(-1/3)_{((\mathcal{K}_{6})^{*},5/6)}).$
Consider smooth curve $C$ and $(f,g)\in(\overline{F}(C)-\overline{F})\times H^{1}(\overline{F}(C),g)$.
$\gamma(f,g)(O)=\frac{1}{3}\cdot\bigg({\deg(f)}{+4\cdot\bigg(\sum_{*\in\{+,-\}}n_{*,1,0}(f,g)+n_{*,4,1}(f,g)\bigg)}+2\cdot\bigg(\sum_{*\in\{+,-\}}n_{*,2,0}(f,g)+n_{*,5,1}(f,g)\bigg)\\ +\bigg(\sum_{*\in\{+,-\}}n_{*,4,0}(f,g)+n_{*,1,1}(f,g)\bigg)-\bigg(\sum_{*\in\{+,-\}}n_{*,5,0}(f,g)+n_{*,2,1}(f,g)\bigg)\bigg)$
Note that $\deg(f)\geq 5n_{+,5,0}(f,g)+2n_{+,2,1}(f,g)$ and that $\deg(f)\geq 5n_{-,5,0}(f,g)+2n_{-,2,1}(f,g)$, hence
$2\deg(f)\geq 5n_{+,5,0}(f,g)+5n_{-,5,0}(f,g)+2n_{+,2,1}(f,g)+2n_{-,2,1}(f,g).$
We deduce $\gamma(f,g)(O)\geq 0.$ In other words, the point $O$ is contained in $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(7))$. If $\gamma(f,g)(O)=0$, for $\in\{+,-\}$, we must have that
$n_{*,5,0}(f,g)=n_{*,4,0}(f,g)=n_{*,1,1}(f,g)=n_{*,2,0}(f,g)=n_{*,5,1}(f,g)=n_{*,1,0}(f,g)=n_{*,4,1}(f,g)=0.$
We must have that $\deg(f)=n_{+,2,1}(f,g)+n_{-,2,1}(f,g)$ and thus, as $\deg(f)\geq 2\max_{*\in\{+,-\}}\{n_{*,2,1}(f,g)\}$, that $n_{+,2,1}(f,g)=n_{-,2,1}(f,g)$. In particular, it follows that all $\gamma(f,g)$ for $(f,g)$ as above are of the form
$(L,x)\mapsto 2mL+mx_{((\mathcal{K}_{6})_{-},5/6)}+mx_{((\mathcal{K}_{6})_{+},5/6)}$
if $\omega\in F$ and
$(L,x)\mapsto 2mL+2mx_{((\mathcal{K}_{6}),1/6)}$
if $\omega\in F$. Let us show that there exists at least one such form $\gamma(f,g)$ for $(f,g)$ as above, not in the span of $(L,x)\mapsto x_{1/2}$. Let’s take $f$ to be $f:\mathbb{P}^{1}\to X_{0}(7)\cong\mathbb{P}^{1}$, given by
$f(z)=\dfrac{\omega^{2}\cdot\big(\frac{z}{z-1}\big)^{2}+\omega}{\big(\frac{z}{z-1}\big)^{2}+1}.$
Note that $f=M\circ h^{2}$, where h(z)=z/(z1)h(z)=z/(z-1) is of degree $1$ and $M(t)=\frac{\omega^{2}t+\omega}{t+1}$ is a Möbius transformation. It follows that $\deg(f)=2$. Moreover, if $f(z)=\omega$, by using that $M(0)=\omega$, we have that $h(z)^{2}=0$, that is $z=0$. Similarly, if $f(z)=\omega^{2}$, by using that $M(\infty)=\omega^{2}$, we obtain $h(z)=\infty$, that is $z=1$. Choose for $g$ any degree $2$ cover of $\mathbb{P}^{1}$, ramified at both $z=0$ and $z=1$. Then
$\gamma(f,g)(L,x)=2L+x_{((\mathcal{K}_{6})_{-},5/6)}+x_{((\mathcal{K}_{6})_{+},5/6)}$
if $\omega\in F$ and
$\gamma(f,g)(L,x)=2L+2x_{((\mathcal{K}_{6}),5/6)}$
if $\omega\not\in F$. By recalling that $(L,x)\mapsto x_{1/2}$ is a form which vanishes on $O$ and which is in $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}$, we obtain that
$\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}\cap\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}|\hskip 2.84544pt\lambda(x)=0\}=\\ =\{(L,x)\mapsto t_{1}x_{1/2}+2t_{2}L+t_{2}x_{((\mathcal{K}_{6})_{-},5/6)}+t_{2}x_{((\mathcal{K}_{6})_{+},5/6)}|\hskip 2.84544ptt_{1},t_{2}\geq 0\}$
if $\omega\in F$ and
$\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}\cap\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}|\hskip 2.84544pt\lambda(x)=0\}=\\ =\{(L,x)\mapsto t_{1}x_{1/2}+2t_{2}L+2t_{2}x_{((\mathcal{K}_{6}),5/6)}|\hskip 2.84544ptt_{1},t_{2}\geq 0\}$
if $\omega\in F$. We deduce that
$b=\dim(\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}\cap\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(7))^{*}|\hskip 2.84544pt\lambda(O)=0\})=2.$
4.5. Case $N=10,25$
Denote by $i$ a chosen square root of 1-1.
Lemma 4.22.
Let $E_{1728}$ be an elliptic curve over $F$ satisfying that $j(E_{1728})=1728$. One has that
$\operatorname{End}(E_{1728})=\begin{cases}\mathbb{Z},&\text{if }i\not\in F\\ \mathbb{Z}[i],&\text{ otherwise}\end{cases}$
Proof.
One has that $\operatorname{End}((E_{1728})_{\overline{F}})=\mathbb{Z}[i]$, hence $\mathbb{Z}\subset\operatorname{End}(E_{0})\subset\mathbb{Z}[i]$. The absolute Galois group $\operatorname{Gal}(\overline{F}/F)$ acts on $\operatorname{End}((E_{1728})_{\overline{F}})$ by $\gamma\cdot\theta=\gamma\circ\theta\circ\gamma^{-1}$. If $\theta$ is represented by $a+b\omega$ with $a,b\in\mathbb{Z}$, the element $\gamma\cdot\theta$ corresponds to
$\gamma\cdot(a+bi)=a+b\gamma\cdot i=a+b(-i)=a-bi$
if $i\not\in F$ and to
$\gamma\cdot(a+bi)=a+bi$
if $i\in F$. We have that $\operatorname{End}(E_{1728})=\operatorname{End}((E_{1728})_{\overline{F}})^{\operatorname{Gal}(\overline{F}/F)}$ which is thus $\mathbb{Z}$ iff $i\in F$ and $\mathbb{Z}[i]$ otherwise. The claim is proven. ∎
Lemma 4.23.
There exist precisely 3[F(i):F]3-[F(i):F] closed points of $\mathscr{X}_{0}(10)$ where the automorphism group is $\mu_{4}$ (they are necessarily lying above the point $j=1728$ in $\mathscr{X}_{0}(1)$ for the morphism $J_{N}:\mathscr{X}_{0}(10)\to\mathscr{X}_{0}(1)$).
Proof.
Suppose that $i\in F$. Then as $\operatorname{End}(E_{1728})=\mathbb{Z}[i]$, we deduce that every cyclic $10$-isogeny of $E_{1728}$ is defined over $F$. By Lemma 3.7, we have that this number is $\varepsilon_{2}(10)=2$. Suppose that $i\not\in F$. Then there is no ideal $\mathfrak{a}$ of $\mathbb{Z}$ satisfying that $N_{\mathbb{Q}(i)/\mathbb{Q}}(\mathfrak{a})=7$, thus there is no isogeny of degree $10$ defined over $F$. It follows that the number of closed points in this case is $\varepsilon_{2}(10)/2=1$. The proof is completed. ∎
Theorem 4.24.
Consider the closed point $K:X^{2}+Y^{2}=0$ of $\mathbb{P}^{1}$. One has that $\mathscr{X}_{0}(10)^{\operatorname{rig}}\cong\sqrt[2]{K}$ and that $\mathscr{X}_{0}(10)$ is a non-trivial root stack $\mathscr{X}_{0}(10)\cong\sqrt[2]{(J_{10}^{\operatorname{rig}})^{*}\Lambda^{\operatorname{rig}}}$.
Proof.
It follows from Lemma 4.23 that the rigidification is $\mathscr{X}_{0}(10)^{\operatorname{rig}}\cong\sqrt[2]{K}$. Let us verify that $\mathscr{X}_{0}(10)\cong\sqrt[2]{J_{10}^{*}(\Lambda^{\operatorname{rig}}})$ is a non-trivial root stack. If we had $\mathscr{X}_{0}(10)\cong\sqrt[2]{K}\times B\mu_{2},$ then $\mathscr{X}_{0}(10)$ would have a point with automorphism group scheme $\mu_{2}\times\mu_{2}$. But we know that such automorphism group schemes do not occur for points in $\mathscr{X}_{0}(10)$. ∎
4.5.1.
Before continuing with the calculation of $a$- and $b$-invariants, let us show that $\mathscr{X}_{0}(25)\cong\mathscr{X}_{0}(10)$. In fact, both stacks are isomorphic to $\mathscr{X}_{0}(5)$.
Lemma 4.25.
If $i\not\in F$, no elliptic curve $E_{1728}$ with $j(E_{1728})=1728$ admits a cyclic $25$-isogeny. If $i\in F$, there exist precisely two pairs consisting of an elliptic curve with $j$ invariant equal to $1728$ and a cyclic $25$-isogeny, which are not isomorphic over $\overline{F}$ and such that any other pair consisting of an elliptic curve with $j$ invariant equal to $1728$ and a cyclic $25$-isogeny is isomorphic to one of them.
Proof.
Suppose that $i\not\in F$. Then $\operatorname{End}(E_{1728})=\mathbb{Z}$. Clearly, the only ideal $\mathfrak{a}$ with the norm $N_{\mathbb{Q}(i)/\mathbb{Q}}(\mathfrak{a})=25$ in $\operatorname{End}(E_{1728})$ is $(5)$. It corresponds to the non-cyclic isogeny given by multiplication by $5$ on $E_{0}$. Suppose that $i\in F$. Then $\operatorname{End}(E_{1728})=\mathbb{Z}[i]$ and every cyclic $25$-isogeny is defined over $F$. But as $\varepsilon_{2}(25)=2$, the second claim follows. ∎
Corollary 4.26.
The stacks $\mathscr{X}_{0}(5)$, $\mathscr{X}_{0}(10)$ and $\mathscr{X}_{0}(25)$ are isomorphic.
Proof.
Combining the proof of Theorem 3.19, Lemma 4.23 and Lemma 4.25 we see that the rigidifications of the three stacks are isomorphic to the root stack $\sqrt{K_{4}}$ where $K_{4}:X^{2}+Y^{2}=0$. Applying the analogous arguments to the ones in the proof of Lemma 4.19, we obtain that $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\cong\mathbb{Z}$ if $i\not\in F$ and $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\cong\mathbb{Z}$ if $i\in F$. We have a diagram $\mathscr{X}_{0}(25)\to\mathscr{X}_{0}(5)\to\mathscr{X}_{0}(1)$ which induces a diagram $\mathscr{X}_{0}(25)^{\operatorname{rig}}\to\mathscr{X}_{0}(5)^{\operatorname{rig}}\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. The composite morphism is of degree $30$, while the right homomorphism is of degree $6$. It follows that the morphism $\mathscr{X}_{0}(25)^{\operatorname{rig}}\to\mathscr{X}_{0}(5)^{\operatorname{rig}}$ is of degree $5$ and similarly, the morphism $\mathscr{X}_{0}(10)^{\operatorname{rig}}\to\mathscr{X}_{0}(5)^{\operatorname{rig}}$ is of degree $3$. Assume $i\not\in F$. Then the induced homomorphism
$\mathbb{Z}/2\mathbb{Z}\cong\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\to\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})\cong\mathbb{Z}/2\mathbb{Z}$
is an isomorphism and similarly
$\mathbb{Z}/2\mathbb{Z}\cong\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\to\operatorname{Pic}(\mathscr{X}_{0}(10)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})\cong\mathbb{Z}/2\mathbb{Z}.$
Hence $(J^{\operatorname{rig}}_{10})^{*}(\Lambda^{\operatorname{rig}})(J^{\operatorname{rig}}_{25})^{*}(\Lambda^{\operatorname{rig}})^{-1}\in 2\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})$ and in this case the claim follows. Assume that $i\in F$. An isomorphism $\mathbb{Z}\oplus(\mathbb{Z}/2\mathbb{Z})\xrightarrow{\sim}\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})$ is given by
$(m,n)\mapsto(J_{5}^{\operatorname{rig}})^{*}(\Lambda)^{\otimes m}\otimes\mathcal{O}((K_{4})_{+}-(K_{4})_{-})^{\otimes n}$
for $m\in\mathbb{Z}$ and $n\in\mathbb{Z}/2\mathbb{Z}$. (As $\mathcal{O}((K_{4})_{+}-(K_{4})_{-})^{\otimes 2}=\mathcal{O}(2(K_{4})_{+}-2(K_{4})_{-})=\pi^{*}\mathcal{O}((K_{4})_{+}-(K_{4})_{-})=0$ where $\pi:\mathscr{X}_{0}(5)^{\operatorname{rig}}\to X_{0}(5)$ is the coarse moduli morphism, the notation is unambiguous.) We will show that the image of the class of the only non-trivial torsion line bundle of $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})$ for $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\to\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})$ is the class of the only non-trivial line bundle in $\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})$. Obviously, it suffices to show that the only non-trivial torsion line bundle for $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\to\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})$ is the only non-trivial torsion line bundle. It follows from Lemma 3.7 that the morphism $\mathscr{X}_{0}(25)^{\operatorname{rig}}\to\mathscr{X}_{0}(5)^{\operatorname{rig}}$ is étale at $K_{4}$. Hence the image of $\mathcal{O}((K_{4})_{+}-(K_{4})_{-})$ is the line bundle $\mathcal{O}(5(K_{4})_{+}-5(K_{4})_{-})$, which is the only non-trivial torsion line bundle. Moreover, it is clear that the image of the class of $(J_{5}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ for $\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(5)^{\operatorname{rig}})\to\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})/2\operatorname{Pic}(\mathscr{X}_{0}(25)^{\operatorname{rig}})$ is the class of $(J_{25}^{*})^{\operatorname{rig}}(\Lambda^{\operatorname{rig}})$. Hence we have an isomorphisms $\mathscr{X}_{0}(5)\xrightarrow{\sim}\mathscr{X}_{0}(25)$ and similarly one shows $\mathscr{X}_{0}(5)\xrightarrow{\sim}\mathscr{X}_{0}(25)$. ∎
4.5.2.
We will fix an implicit identification $X_{0}(N)\cong\mathbb{P}^{1}$. Given an $\overline{F}$-point of a curve $x$, we denote by $V_{x}$ the corresponding formal neighborhood. Given a pair $(f,c)\in(\overline{F}(C)-\overline{F})\times H^{1}(\overline{F}(C),\mu_{2})$ with $C$ smooth projective irreducible $\overline{F}$-curve, we define:
  • $n_{1/2}(f,g)$ to be the number of geometric points $x$ of $C$ above which $g$ is ramified and which satisfy that $f(x)\not\in K$;
  • Let $(\mathcal{K}_{4})_{\overline{F}}=(\mathcal{K}_{4})_{+}\cup(\mathcal{K}_{4})_{-}$ be the decomposition of $(\mathcal{K}_{4})_{\overline{F}}$ into two disjoint points. Denote $(K_{4})_{*}$ the image of $(\mathcal{K}_{4})_{*}$ with $\in\{+,-\}$ for the coarse moduli space map and $K_{4}=(K_{4})_{+}\cup(K_{4})_{-}$. We set $n_{+,i,0}(f,g)$ (respectively, $n_{-,1,0}(f,g)$) to be the number of geometric points $x$ of $C$ above which $g$ is unramified, which satisfy that $f(x)=(K_{4})_{+}$ (respectively, $f(x)=(K_{4})_{-}$) and the ramification of $f$ at $x$ is congruent $i$ modulo $4$.
  • We set $n_{+,1,1}(f,g)$ (respectively, $n_{-,1,1}(f,g)$) to be the number of geometric points $x$ of $C$ above which $g$ is ramified, which satisfy that $f(x)=(K_{4})_{+}$ (respectively, $f(x)=(K_{4})_{-}$) and the ramification of $f$ is congruent $i$ modulo $4$.
For a pair $(f,g)$ as above, we denote by $\gamma(f,g)$ the following element of $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(10))^{*}$:
  • Assume that $i\not\in F$. We set:
    $\gamma(f,g)(L,x):=\deg(f)\cdot L+x_{((\mathcal{K}_{4}),1/4)}\cdot\sum_{*\in\{+,-\}}(n_{*,1,0}(f,g)+n_{*,3,1}(f,g))+\\ +x_{1/2}\cdot\bigg(n_{1/2}(f,g)+\sum_{*\in\{+,-\}}n_{*,0,1}(f,g)+n_{*,2,1}(f,g)\bigg)+x_{((\mathcal{K}_{4}),3/4)}\cdot\sum_{*\in\{+,-\}}(n_{*,1,1}(f,g)+n_{*,3,0}(f,g)).$
  • Assume that $i\in F$. We set:
    $\gamma(f,g)(L,x):=\deg(f)\cdot L+\sum_{*\in\{+,-\}}x_{((\mathcal{K}_{4})_{*},1/4)}\cdot(n_{*,3,0}(f,g)+n_{*,1,1})+\\ +x_{1/2}\cdot\bigg(n_{1/2}(f,g)+\sum_{*\in\{+,-\}}n_{*,0,1}(f,g)+n_{*,2,1}(f,g)\bigg)+\sum_{*\in\{+,-\}}x_{((\mathcal{K}_{4})_{*},3/4)}\cdot(n_{*,1,0}(f,g)+n_{*,3,1}(f,g)).$
It follows from Lemma 4.5, that $\gamma(f,g)\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))^{*}$ is induced by $\iota(f,g)^{\nu}$.
Theorem 4.27.
When $N=10$ we have that $a=1/6$ and that $b=1$. When $N=25$, we also have that $a=1/6$ and that $b=1$.
Proof.
Using notation $(\mathcal{K}_{4})_{*}$ as before in the case $i\in F$ and using it for $(\mathcal{K}_{4})$ in the case $i\not\in F$, one has that
$K_{\mathscr{X}_{0}(10),\operatorname{orb}}=(-1/2,(-1/2)_{((\mathcal{K}_{4})_{*},1/4)},(-1)_{1/2},(-1/2)_{(\mathcal{K}_{4})_{*},3/4}).$
When $N=10$, the naive height corresponds to
$T:=(9,(9)_{((\mathcal{K}_{4})_{*},1/4)},(6)_{1/2},(3)_{((\mathcal{K}_{4})_{*},3/4)}).$
We want the infimum of $a$ for which $aT+K_{\mathscr{X}_{0}(10),\operatorname{orb}}\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(10))$. Recall that $x_{1/2}\geq 0$ is an equation defining cone, so $a\geq 1/6$. Consider the point
$O:$ $=(1/6)\cdot T+K_{\mathscr{X}_{0}(10)}$
$=(3/2,(3/2)_{((\mathcal{K}_{4})_{*},1/4)},(1)_{1/2},(1/2)_{(K_{*},3/4)})+(-1/2,(-1/2)_{((\mathcal{K}_{4})_{*},1/4)},(-1)_{1/2},(-1/2)_{K_{*},3/4})$
$=(1,(1)_{((\mathcal{K}_{4})_{*},1/4)},(0)_{1/2},(0)_{(\mathcal{K}_{4})_{*},3/4}).$
Consider $(f,g)$ as before. Assume that $\omega\in F$. We have that
$\gamma(f,g)(O)\geq\deg(f)\cdot 1>0.$
Hence the inequalities defined by $(f,g)$ as above are all strict. We get $a=1/6$. We also get
$b=\dim(\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(10))^{*}\cap\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(10))^{*}|\hskip 2.84544pt\lambda(O)=0\})=1.$
When $N=25$, the naive height corresponds to
$T^{\prime}:=(15,(9)_{(K_{*},1/4)},(6)_{1/2},(3)_{(K_{*},3/4)}).$
We want the infimum of $a$ for which $aT^{\prime}+K_{\mathscr{X}_{0}(10),\operatorname{orb}}\in\operatorname{\overline{Eff}}(\mathscr{X}_{0}(10))$. Again, as $x_{1/2}\geq 0$ is an equation defining cone, we have $a\geq 1/6$. Consider the point
$O^{\prime}:$ $=(1/6)T^{\prime}+K_{\mathscr{X}_{0}(10)}$
$=(5/2,(3/2)_{((\mathcal{K}_{4})_{*},1/4)},(1)_{1/2},(1/2)_{((\mathcal{K}_{4})_{*},3/4)})+(-1/2,(-1/2)_{((\mathcal{K}_{4})_{*},1/4)},(-1)_{1/2},(-1/2)_{(\mathcal{K}_{4})_{*},3/4})$
$=(2,(1)_{((\mathcal{K}_{4})_{*},1/4)},(0)_{1/2},(0)_{(\mathcal{K}_{4})_{*},3/4}).$
Consider $(f:C\to\mathbb{P}^{1},g:\widetilde{C}\to C)$ as before. We have that
$\gamma(f,g)(O^{\prime})\geq\deg(f)\cdot 2>0.$
We obtain
$a=1/6\text{ and }b=\dim(\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(10))^{*}\cap\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(10))^{*}|\hskip 2.84544pt\lambda(O^{\prime})=0\})=1.$
The proof is completed. ∎
4.6. Case $N=13$
Lemma 4.28.
There exist precisely 3[F(i):F]3-[F(i):F] closed points of $\mathscr{X}_{0}(13)$ with automorphism group scheme $\mu_{4}$ (they are necessarily lying above the point $j=1728$ in $\mathscr{X}_{0}(1)$ for the morphism $J_{N}:\mathscr{X}_{0}(13)\to\mathscr{X}_{0}(1)$). There exist precisely $3-[F(\omega):F]$ closed points of $\mathscr{X}_{0}(13)$ with automorphism group scheme $\mu_{6}$ (they are necessarily lying above the point $j=0$ in $\mathscr{X}_{0}(1)$ for the morphism $J_{N}:\mathscr{X}_{0}(13)\to\mathscr{X}_{0}(1)$)
Proof.
An analogous proof as before applies. ∎
As usual we fix an identification $X_{0}(N)\cong\mathbb{P}^{1}$, with the image of $\mathcal{K}_{4}$ and $\mathcal{K}_{6}$ under the coarse moduli space denoted by $K_{4}$ and $K_{6}$. Up to an automorphism we can identify $K_{4}$ and $K_{6}$ with the closed points given in the following lemma.
Theorem 4.29.
Consider the closed points $K_{4}:X^{2}+Y^{2}=0$ and $K_{6}=X^{2}+3Y^{2}$ of $\mathbb{P}^{1}$. One has that $\mathscr{X}_{0}(13)^{\operatorname{rig}}\cong\sqrt[2,3]{K_{4},K_{6}},$ where the notation means that $\mu_{2}$ stackiness is added along $K_{4}$ and $\mu_{3}$ stackiness is added along $K_{6}$. Moreover, one has that $\mathscr{X}_{0}(13)$ is a non-trivial root stack $\mathscr{X}_{0}(13)\cong\sqrt[2]{J_{13}^{*}(\Lambda^{\operatorname{rig}}})$
Proof.
It follows from Lemma 4.28 that the rigidification is $\mathscr{X}_{0}(13)^{\operatorname{rig}}\cong\sqrt[2,3]{K,D}$. An analogous argument to the one for $N=10$ shows that $\mathscr{X}_{0}(13)$ is a non-trivial gerbe $\mu_{2}$-gerbe over $\sqrt[2,3]{K,D}$. ∎
Given a smooth projective $\overline{F}$-curve $C$ and a pair $(f,g)\in(\overline{F}(C)-\overline{F})\times H^{1}(\overline{F}(C),\mu_{2})$. By $((\mathcal{K}_{6})_{*},a)$ we mean $((\mathcal{K}_{6}),a)$ if $\omega\not\in F$ and it has its standard meaning if $\omega\in F$ and by $((\mathcal{K}_{4})_{*},a)$ we mean $((\mathcal{K}_{4}),a)$ if $i\not\in F$ and it has its standard meaning if $i\in F$. We define $w(1,0)=w(4,1)=1/6$, $w(4,0)=w(1,1)=2/3$, $w(3,0)=w(3,1)=1/2$, $w(2,0)=w(5,1)=1/3$, and $w(1,0)=w(5,1)=5/6$. Similarly, we define $q(3,0)=q(1,1)=1/4$, $q(2,0)=q(2,1)=1/2$ and $q(1,0)=q(3,1)=3/4$. Write $n_{*,(3,0)}$ and $n_{*,(3,1)}$ for $n_{1/2}.$ We define:
$\gamma(f,g)(L,x):=\deg(f)\cdot L+\sum_{*\in\{+,-\}}\sum_{j=1}^{5}\sum_{r\in w^{-1}(j/6)}n_{*,r}(f,g)x_{((\mathcal{K}_{6})_{*},j/6)}+\\ +\sum_{*\in\{+,-\}}\sum_{j=1}^{3}\sum_{r\in q^{-1}(j/4)}n_{*,r}(f,g)x_{((\mathcal{K}_{4})_{*},j/4)}.$
Theorem 4.30.
In the case $N=13$, one has that $a=1/6$ and $b=1$.
Proof.
One has that
$K_{\mathscr{X}_{0}(13),\operatorname{orb}}=\big(1/6,(-1/2)_{((\mathcal{K}_{4})_{*},1/4)},(-1/2)_{((\mathcal{K}_{4})_{*},3/4)},\\ (-1)_{1/2},(-2/3)_{((\mathcal{K}_{6})_{*},1/6)},(-1/3)_{((\mathcal{K}_{6})_{*},1/3)},(-2/3)_{((\mathcal{K}_{4})_{*},2/3)},(-1/3)_{((\mathcal{K}_{4})_{*},5/6)}\big).$
The naive height is given by the element
$T:=(7,(9)_{((\mathcal{K}_{4})_{*},1/4)},(3)_{((\mathcal{K}_{4})_{*},3/4)},(6)_{1/2},(10)_{((\mathcal{K}_{6})_{*},1/6)},(8)_{((\mathcal{K}_{6})_{*},1/3)},(4)_{((\mathcal{K}_{6})_{*},2/3)},(2)_{((\mathcal{K}_{6})_{*},5/6)})$
We want the infimum of $a$ for which $aT+K_{\mathscr{X}_{0}(13),\operatorname{orb}}\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(13))$. Recall that $x_{1/2}\geq 0$ is an equation defining cone, so $a\geq 1/6$. Consider the point
$O:$ $=(1/6)\cdot T+K_{\mathscr{X}_{0}(13),\operatorname{orb}}$
$=(4/3,(1)_{((\mathcal{K}_{4})_{*},1/4)},(0)_{((\mathcal{K}_{4})_{*},3/4)},(0)_{1/2},(4/3)_{((\mathcal{K}_{6})_{*},1/6)},(2/3)_{((\mathcal{K}_{6})_{*},1/3)},(1/3)_{((\mathcal{K}_{6})_{*},2/3)},(-1/3)_{((\mathcal{K}_{6})_{*},5/6)}).$
Let $C$ be a smooth projective irreducible $\overline{F}$-curve, and $(f,g)$ as usual. We have that
$\gamma(f,g)(O)\geq 4/3\cdot\deg(f)+(-1/3)\sum_{*\in\{+,-\}}(n_{*,5,0}(f,g)+n_{*,1,1}(f,g)).$
We have that $\deg(f)\geq 5n_{+,5,0}(f,g)+n_{+,1,1}(f,g)$ and $\deg(f)\geq 5n_{-,5,0}(f,g)+n_{-,1,1}(f,g)$. We obtain
$4\deg(f)$ $\geq 10n_{+,5,0}(f,g)+2n_{+,1,1}(f,g)+10n_{n-,5,0}(f,g)+2n_{-,1,1}(f,g)$
$\geq\sum_{*\in\{+,-\}}(n_{*,5,0}(f,g)+n_{*,1,1}(f,g)).$
Hence $\gamma(f,g)(O)\geq 0$ for all $f,g$ as above. Thus $O\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(13))$. Hence $a=1/6.$ For the equality $\gamma(f,g)(O)=0$ to hold, one must have
$n_{*,5,0}(f,g)=n_{*,1,1}(f,g)=0$
for $\in\{+,-\}$. But as $\deg(f)\geq 1$, we have that $\gamma(f,g)(O)>0$, so the equality is not achieved for any $f,g$ as above. It follows that
$\{\lambda\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(13))^{*}|\hskip 2.84544pt\lambda(O)=0,\lambda\in\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(13))^{*}\}=\{(L,x)\mapsto tx_{1/2}|\hskip 2.84544ptt\geq 0\}$
which is of dimension $1$. Hence $b=1$. The claim is proven. ∎